### v6

# Information Loss under Coarse-Grained Partitions: A Discrete Framework for AI Ethics

Abstract
As artificial intelligence (AI) systems become increasingly embedded in ethically sensitive domains such as education, healthcare, and transportation, the need to balance accuracy and interpretability in decision-making has become a central concern. Coarse Ethics (CE) motivates the use of coarse-grained evaluations, such as letter grades, as ethically appropriate under cognitive, institutional, and contextual constraints. However, CE has lacked a simple mathematical formalization that makes the informational cost of such coarse judgments explicit. This paper introduces coarse-grained partitions (CGPs) as a discrete, information-theoretic framework for modeling coarse evaluation. We treat an underlying score scale as a finite ordered set, represent coarse evaluation as a partition into grains together with an index assignment, and define the induced coarse-grained distribution by pushforward. To quantify the informational impact of coarse-graining, we propose categorical unification (CU), which constructs a canonical within-grain unification of the original distribution, and we measure information loss by the discrete Kullback–Leibler divergence between the original distribution and its CU-unification (KL-CU). We show that information loss is zero if and only if the original distribution is already uniform within each grain, clarifying precisely when coarse evaluation preserves all relevant information. We illustrate the framework with educational grading and explainable AI (XAI) settings, where coarse categories are often required for communication and accountability. By providing a tractable discrete formalization of coarse evaluation and a principled measure of information loss, CGPs offer a foundation for designing interpretable AI systems that balance comprehensibility, fairness, and informational integrity.
keywords:
Coarse-Grained Partitions, AI Ethics, Coarse Ethics, Explainable AI, Coarse-Grained AI Models, KL Divergence
1 Introduction
1.1 Background
As artificial intelligence (AI) becomes increasingly embedded in decision-making processes, the need for effective user interaction with AI-driven systems has become a critical research concern. Concepts such as fairness, transparency, and explainability, including explainable AI (XAI), have been central to this discourse [gunning2019xai]. However, a well-established trade-off exists between predictive accuracy and interpretability in AI models, where higher precision often results in reduced explainability [BARREDOARRIETA202082].
To address this challenge, researchers have explored various approximation techniques, which can be broadly categorized into engineering-driven and ethics-driven approaches. From an engineering perspective, a common strategy involves approximating complex black-box models with more interpretable alternatives. Techniques such as Local Interpretable Model-Agnostic Explanations (LIME, [ribeiro2016whyitrustyou]) and SHapley Additive exPlanations (SHAP, [lundberg2017unifiedapproachinterpretingmodel]) have been widely used to provide post-hoc interpretability for AI systems.
From an ethical standpoint, Coarse Ethics (CE, [izumoweng2022]) proposes that coarse-grained assessments are not only necessary but also ethically justifiable in many decision-making contexts. For example, academic grading systems classify students into letter grades or grade point average (GPA), abstracting away precise score differences. Similarly, phrases like “Olivia and Noah are Harvard graduates” generalize individual performance details. In AI-driven decision-making, CE argues that the legitimacy of transforming fine-grained evaluations into coarse ones depends on the cognitive capacity and contextual needs of the audience [izumo2025].
1.2 Challenges in Coarse Ethics
Despite its conceptual appeal, CE faces two primary challenges:
  1. 1.
    Lack of Mathematical Rigor: While CE provides an ethical justification for coarse evaluations, it lacks a mathematically formalized framework. For instance, Izumo [izumo2025] proposes that an AI assistant should issue coarse warnings such as “Your driving is dangerous” rather than continuously recalculating and displaying the optimal speed. However, the study does not specify the exact computational method to determine when and how such warnings should be issued. Without a formalized structure, these decisions may be inconsistent, raising concerns about reliability in AI-driven risk assessment.
  2. 2.
    Ambiguity in Categorization: Prior research has introduced a three-tiered explanatory approach—detailed explanations for experts, coarse explanations for general users, and even coarser explanations for individuals with cognitive limitations [izumo2025]. However, this classification is overly simplistic and lacks adaptability, failing to account for the diverse range of cognitive and contextual factors in human decision-making.
Addressing these limitations is essential, particularly as AI advances toward artificial general intelligence (AGI). Strengthening the mathematical foundation of CE and refining its categorization mechanisms can contribute to the development of more robust, context-aware AI explanations.
1.3 Methodology
Several prior studies share conceptual similarities with CE. For example, He et al. [he2024analogies] explored how analogies can assist laypeople in AI-assisted decision-making. Their work emphasizes the cognitive aspects of explanation reception and user comprehension. While our study shares the objective of improving human-AI interaction, we focus on the mathematical quantification of information loss in coarse evaluations rather than the psychological mechanisms of explainability. Peters and Carman [peters2024cultural] examined the influence of cultural bias in XAI research, demonstrating how users from different cultural backgrounds interpret AI explanations differently. Their study focuses on the subjective perception of AI explanations rather than developing a mathematical framework for evaluation. In contrast, our work formalizes coarse-grained evaluations using set theory, providing a structured approach to managing granularity in AI-driven decision-making.
This paper proposes coarse-grained partitions (CGPs) as a novel set-theoretic approach to resolving these issues. Set theory has been extensively applied to human decision-making, forming the basis of fuzzy set theory [zadeh1965fuzzy], rough set theory [pawlak1982rough], and soft set theory [molodtsov1999soft]. These approaches primarily focus on modeling real-world uncertainty through mathematical representations.
In contrast, CGPs shift the focus from enhancing AI’s ability to analyze the world to helping humans extract and interpret AI-generated insights. While previous works aimed at improving AI’s analytical capabilities with set theory [10.5120/8924-2996, 9015977, dalkilic2022novel], CGPs are based on the premise that AI models now outperform human experts in specific domains, as demonstrated by recent studies on AI surpassing human-level performance in standardized exams [bicknell2024chatgpt, katz2023gpt4]. Consequently, the key question is no longer how AI perceives reality, but rather how humans can efficiently understand and interact with AI-driven decisions.
This perspective is particularly relevant as Large Language Models (LLMs) and domain-specific AI systems continue to exceed human performance in complex tasks [vaswani2017attention, radford2018improving, brown2020language, chowdhery2022palm]. As AI models continue to outperform human experts in specialized tasks, discussions surrounding AGI expectations and ethical considerations have intensified [morris2024levelsagioperationalizingprogress]. By establishing a formal mathematical foundation for coarse evaluations, CGPs offer a critical framework for addressing these challenges and redefining the evolving relationship between AI and human cognition.
2 Coarse-Grained Partitions
Throughout the paper, we fix a finite totally ordered set $(U,\leq)$, which we call the underlying scale. In applications, one may take $U=\{0,1,\dots,100\}$ with the natural order. We equip $U$ with the discrete $\sigma$-algebra $2^{U}$.
2.1 Coarse-Grained Partitions
Definition 1 (Coarse-grained partition).
Let $(U,\leq)$ be a finite totally ordered set, and let $\mathscr{I}(U,\leq)$ denote the family of all nonempty order-convex (interval) subsets of $U$. A coarse-grained partition of $(U,\leq)$ is a partition object $\pi$ equipped with a finite index set $I_{\pi}$ and a family of grains
$\mathfrak{G}_{\pi}:=\{G_{\pi,i}\}_{i\in I_{\pi}}$
such that: (i) each $G_{\pi,i}\in\mathscr{I}(U,\leq)$ is a nonempty interval; (ii) $G_{\pi,i}\cap G_{\pi,j}=\varnothing$ for $i\neq j$; (iii) $\bigcup_{i\in I_{\pi}}G_{\pi,i}=U$. We call each $G_{\pi,i}$ a grain (or category) of $\pi$.
Here $\pi$ denotes the partition object, while $\mathfrak{G}_{\pi}$ denotes its underlying family of grains.
Remark 1 (On generality).
The discrete ordered framework above is sufficient for the main results of this paper and matches typical applications such as grading scales. A measure-theoretic extension to general measurable spaces can be developed separately (e.g. by replacing sums with integrals), and may be placed in an appendix if needed.
Remark 2 (Ordering grain families).
A key structural feature of interval coarse-graining is that it induces an order not only on $U$ but also on the grain family itself. Define, for $G,H\in\mathfrak{G}_{\pi}$,
$G\preccurlyeq H\quad:\Longleftrightarrow\quad(\forall x\in G)(\forall y\in H)\;x\leq y.$
This element-wise relation captures a global ordering of categories and is compatible with the intuitive notion that some coarse-grained categories are uniformly “lower” (or “worse”) than others along the evaluation scale. Although the present paper focuses on probabilistic pushforwards and information loss under within-grain unification, this induced grain order provides the missing structure needed to define order-sensitive operators on partitions (e.g., adjacent merging, monotone coarsening, and order-constrained optimization of coarse-grainings).
Remark 3 (No internal order within grains).
In our constructions, each grain $G\in\mathfrak{G}_{\pi}$ is treated as an unordered set: only membership $u\in G$ and the cardinality $|G|$ are used. In particular, the CU construction assigns probabilities uniformly within each grain, so no internal ordering of elements inside a grain is required in this paper. By contrast, in extensions where one selects a representative value for each grain—for instance, taking $\min(G)$ as the representative—the ambient order on $U$ (and hence the induced order within $G$) becomes relevant.
Example 1.
Let $U_{1}:=\{1,2,3\}$ be a finite totally ordered set with the natural order $\leq$. The nonempty intervals (order-convex subsets) of $(U_{1},\leq)$ are
$\{1\},\ \{2\},\ \{3\},\ \{1,2\},\ \{2,3\},\ \{1,2,3\}.$
Equivalently,
$\mathscr{I}(U_{1},\leq)=\bigl\{\{1\},\{2\},\{3\},\{1,2\},\{2,3\},\{1,2,3\}\bigr\}.$
Applying Definition 1, there are exactly four interval coarse-grained partitions of $U_{1}$. Writing each partition $\pi_{k}$ through its grain family $\mathfrak{G}_{\pi_{k}}$, we obtain:
$\mathfrak{G}_{\pi_{1}}$ $=\bigl\{\{1\},\{2\},\{3\}\bigr\},$
$\mathfrak{G}_{\pi_{2}}$ $=\bigl\{\{1,2\},\{3\}\bigr\},$
$\mathfrak{G}_{\pi_{3}}$ $=\bigl\{\{1\},\{2,3\}\bigr\},$
$\mathfrak{G}_{\pi_{4}}$ $=\bigl\{\{1,2,3\}\bigr\}.$
In this context, each element of the grain family associated with a partition (e.g., $\{1,2\}$ in $\mathfrak{G}_{\pi_{2}}$) is a grain (interval) of $U_{1}$, so this terminology is natural.
The collection of (interval) coarse-grained partitions of $U$ is denoted by $\mathcal{P}_{\textrm{int}}(U)$ in this paper. In particular, for the example $U_{1}=\{1,2,3\}$ above, we have
$\mathcal{P}_{\textrm{int}}(U_{1})=\{\pi_{1},\pi_{2},\pi_{3},\pi_{4}\},$
where each partition $\pi_{k}$ is represented by its grain family $\mathfrak{G}_{\pi_{k}}$ as listed in the example. Accordingly, the collection of grain families induced by interval partitions is
$\mathscr{G}_{\textrm{int}}(U_{1})=\{\mathfrak{G}_{\pi_{1}},\mathfrak{G}_{\pi_{2}},\mathfrak{G}_{\pi_{3}},\mathfrak{G}_{\pi_{4}}\}.$
If arbitrary set partitions were allowed, then the number of partitions of an $n$-element set would be the Bell number $B_{n}$, which grows extremely rapidly. By contrast, restricting to order-convex (interval) grains makes each partition equivalent to choosing cut points between adjacent elements; hence there are exactly $2^{\,n-1}$ interval partitions of a finite totally ordered $n$-element set. This restriction rules out nonlocal groupings while keeping the combinatorial complexity of the partition space substantially lower than in the unrestricted case.
Proposition 1 (Counting interval partitions).
Let $U$ be a finite totally ordered set with $|U|=n$. Then the number of coarse-grained partitions of $U$ into order-convex grains (i.e., interval partitions) is
$2^{\,n-1}.$
Proof of Proposition 1.
List the elements of $U$ in increasing order as $u_{1}\prec u_{2}\prec\cdots\prec u_{n}$. An interval partition is uniquely determined by deciding, for each of the n1n-1 gaps between consecutive elements $(u_{i},u_{i+1})$, whether to place a cut (start a new grain) or not. Each gap offers two choices, independently. Hence there are $2^{\,n-1}$ interval partitions. ∎
3 Coarse Assignment and Category Assignment
3.1 Coarse Assignment
CE studies how a fine-grained evaluation on a scale $U$ is mapped into a coarser categorization. Given a coarse-grained partition $\pi$ of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$, the coarse categories are conceptually given by the grains $G_{\pi,i}\in\mathfrak{G}_{\pi}$. For technical convenience (measurability and pushforward constructions), we represent these categories by their labels $I_{\pi}$ and fix a coarse assignment map
$q_{\pi}:U\to I_{\pi},\qquad q_{\pi}(x)=i\ \Longleftrightarrow\ x\in G_{\pi,i}.$
Definition 2 (Index set and coarse assignment map).
Let $(U,\leq)$ be a finite totally ordered set, equipped with the discrete $\sigma$-algebra $2^{U}$. Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ (cf. Def. 1). Define the index set by
$I_{\pi},$
equipped with the discrete $\sigma$-algebra $2^{I_{\pi}}$. The coarse assignment map associated with $\pi$ is the map $q_{\pi}:U\to I_{\pi}$ defined by
$q_{\pi}(x)=i\quad\Longleftrightarrow\quad x\in G_{\pi,i}.$
Example 2 (A coarse assignment map).
Let $U=\{1,2,3\}$ with the discrete $\sigma$-algebra $2^{U}$. Define a coarse-grained partition $\pi$ of $U$ by the grain family
$\mathfrak{G}_{\pi}:=\{G_{\pi,i}\}_{i\in I_{\pi}}=\bigl\{\{1\},\{2,3\}\bigr\},\qquad I_{\pi}:=\{1,2\}.$
Define the coarse assignment map $q_{\pi}:U\to I_{\pi}$ by $q_{\pi}(u)=i$ iff $u\in G_{\pi,i}$. Then
$q_{\pi}(1)=1,\qquad q_{\pi}(2)=2,\qquad q_{\pi}(3)=2.$
Remark 4 (Measurability in the finite/discrete setting).
Since we work with the discrete $\sigma$-algebras $2^{U}$ and $2^{I_{\pi}}$ on finite sets, every map between these spaces is measurable. In particular, the coarse assignment map $q_{\pi}:U\to I_{\pi}$ is automatically measurable.
3.2 Object-to-Category Map
In practice, an evaluation is always performed of some collection of objects. Therefore, it is not enough to specify only the relationship between a fine-grained scale and its coarse-grained counterpart. One must also fix (i) how objects are mapped to fine-grained scores, and (ii) how those scores are subsequently mapped to coarse categories. In other words, we must specify both the object-to-score map and the induced object-to-category assignment.
Definition 3 (Object-to-category map).
Let $O$ be a nonempty set of objects and let $f:O\to U$ be an object-to-score map. Fix a coarse-grained partition $\pi$ of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ and coarse assignment map $q_{\pi}:U\to I_{\pi}$. The object-to-category map induced by $f$ and $\pi$ is the composition
$q_{\pi}\circ f:O\to I_{\pi},\qquad o\mapsto q_{\pi}(f(o)).$
3.3 Illustration by School Test
Let’s consider a specific example. Suppose Olivia, Noah, and James took a mathematics test and scored out of 100 points. The scores are then converted into letter grades as follows: 0–59 points correspond to poor, 60–69 to fair, 70–79 to good, 80–89 to very good, and 90–100 to excellent. In this evaluation system, a scale from 0 to 100 is the underlying set $U$, and the letter grades represent its grains.
Let
$U:=\{0,1,2,\dots,99,100\}\quad\text{with the natural order }\leq\text{ on }U.$
Let $I_{\pi}:=\{1,2,3,4,5\}$ and define a finite coarse-grained partition $\pi$ of $U$ by
$\mathfrak{G}_{\pi}:=\{G_{\pi,i}\}_{i\in I_{\pi}},$
where the grains are given by
$G_{\pi,1}$ $:=\{0,\dots,59\},$ $G_{\pi,2}$ $:=\{60,\dots,69\},$
$G_{\pi,3}$ $:=\{70,\dots,79\},$ $G_{\pi,4}$ $:=\{80,\dots,89\},$
$G_{\pi,5}$ $:=\{90,\dots,100\}.$
For readability, introduce a label map $\ell:I_{\pi}\to L$ where
$L:=$ $\{\text{poor},\text{fair},\text{good},\text{very good},\text{excellent}\},$
$\ell(1)=\text{poor},\ \ell(2)=\text{fair},\ \ell(3)=\text{good},\ \ell(4)=\text{very good},\ \ell(5)=\text{excellent}.$
Then:
$O=\{\text{Olivia},\text{Noah},\text{James}\},$
$f:O\to U;\quad f(\text{Olivia})=90,\ f(\text{Noah})=71,\ f(\text{James})=77,$
$q_{\pi}:U\to I_{\pi};\quad q_{\pi}(x)=i\Longleftrightarrow x\in G_{\pi,i},$
$\ell\circ q_{\pi}\circ f:O\to L;$
$\quad(\ell\circ q_{\pi}\circ f)(\text{Olivia})=\text{excellent},\ (\ell\circ q_{\pi}\circ f)(\text{Noah})=\text{good},\ (\ell\circ q_{\pi}\circ f)(\text{James})=\text{good}.$
In this case, comparing the mappings $f:O\to U$ and $\ell\circ q_{\pi}\circ f:O\to L$ is informative. While $f$ distinguishes all three individuals by their exact scores, the coarse evaluation collapses distinct scores (e.g., Noah and James) into the same label “good”. This loss of within-grain distinction affects the induced frequencies over labels, as discussed in the next section.
Remark 5.
The role of the function $q_{\pi}$ can be understood as follows: if the three students are aware only of the grading system and the results, they perceive their results as multivalued information. For example, if they are told that Olivia received an “excellent” grade, they know that her score was one of 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, or 100. In other words, the role of $q_{\pi}$ is to replace the information that Olivia scored 90 with the information that her score was one of 90 through 100. This means that the original information is lost, which gives practical value to the considerations in Section 5.
4 Information Loss
4.1 Probability Measures
Since normative assessments are typically made over a finite set of cases in practice, this section restricts attention to the finite setting. We model both objects and scores as finite sets and represent empirical frequencies by discrete probability distributions. This choice keeps the information-loss analysis entirely within the standard discrete (Shannon/KL) formalism, without comparing probability measures of different types (e.g., continuous vs. atomic).
Definition 4 (Pushforward (coarse-grained) probability).
Let $U$ be a finite set and equip it with the power set $\Sigma:=2^{U}$. Let $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ be a finite coarse-grained partition of $U$, and let $P_{U}$ be a probability distribution on $(U,2^{U})$. With $q_{\pi}$ as in Definition 2, define the coarse-grained probability distribution $P_{\pi}$ on $(I_{\pi},2^{I_{\pi}})$ by the pushforward
$P_{\pi}:=P_{U}\circ q_{\pi}^{-1}.$
Equivalently, for each $i\in I_{\pi}$,
$P_{\pi}(\{i\})=P_{U}(G_{\pi,i}).$
Remark 6 (The coarse $\sigma$-algebra generated by $\pi$).
The partition $\pi$ induces the $\sigma$-algebra
$\Sigma_{\pi}:=\sigma(\pi)=\Bigl\{\bigcup_{i\in J}G_{\pi,i}\;:\;J\subseteq I_{\pi}\Bigr\}\subseteq\Sigma.$
Moreover, $q_{\pi}$ is $(\Sigma,\;2^{I_{\pi}})$-measurable and $\Sigma_{\pi}=q_{\pi}^{-1}(2^{I_{\pi}})$.
The space $O$ represents the underlying domain of objects (e.g., test scores), equipped with a probability distribution $\mathbb{P}$ that captures how likely different objects are to occur. The map $f:O\to U$ assigns each object in $O$ to a category in $U$. Definition 4 defines the induced probability distribution on $U$ by pushing forward $\mathbb{P}$ through $f$: for any measurable set of categories $A\subseteq U$, the probability that the assigned category lies in $A$ is defined as the probability mass of all objects in $O$ that are mapped into $A$, i.e.,
$\mathbb{P}_{U}(A)=\mathbb{P}(f^{-1}(A)).$
In other words, $\mathbb{P}_{U}$ is the distribution of the categorized outcome $f(o)$ when $o$ is drawn according to $\mathbb{P}$.
Example 3 (Pushforward to a win/lose outcome).
As a running toy example, consider a dice game in which a larger face value yields a higher score. We take the underlying scale to be
$U:=\{1,2,3,4,5,6\}\quad\text{equipped with the natural total order }1\leq 2\leq\cdots\leq 6.$
Let $O:=\{a,b,c,d,e\}$ be a set of five players (or five trials), and define an outcome map $f:O\to U$ by
$f(a)=1,\quad f(b)=1,\quad f(c)=3,\quad f(d)=5,\quad f(e)=6.$
The induced empirical distribution $\hat{P}\in\Delta(U)$ is defined by
$\hat{P}(u):=\frac{|\{o\in O:f(o)=u\}|}{|O|}\qquad(u\in U).$
Hence,
$\hat{P}(1)=\frac{2}{5},\ \hat{P}(2)=0,\ \hat{P}(3)=\frac{1}{5},\ \hat{P}(4)=0,\ \hat{P}(5)=\frac{1}{5},\ \hat{P}(6)=\frac{1}{5}.$
Next, define a two-grain partition $\pi$ of $U=\{1,2,3,4,5,6\}$ by the grain family
$G_{\pi,1}:=\{1,2,3\},\qquad G_{\pi,2}:=\{4,5,6\},$
and denote the associated grain family by
$\mathfrak{G}_{\pi}:=\{G_{\pi,1},G_{\pi,2}\}.$
(For readability, we may refer to $G_{\pi,1}$ as the Lose-grain and $G_{\pi,2}$ as the Win-grain.)
Let $I_{\pi}:=\{1,2\}$ and define the coarse assignment map $q_{\pi}:U\to I_{\pi}$ by
$q_{\pi}(u)=\begin{cases}1,&u\in G_{\pi,1}=\{1,2,3\},\\ 2,&u\in G_{\pi,2}=\{4,5,6\}.\end{cases}$
Introduce the grain-label map $\ell:I_{\pi}\to\mathfrak{G}_{\pi}$ by
$\ell(1)=G_{\pi,1},\qquad\ell(2)=G_{\pi,2}.$
Then $\ell\circ q_{\pi}:U\to\mathfrak{G}_{\pi}$ maps each outcome $u\in U$ to its grain, and hence
$\ell\circ q_{\pi}\circ f:O\to\mathfrak{G}_{\pi}.$
The pushforward distribution on grains, $\hat{P}_{\mathfrak{G}}:=(\ell\circ q_{\pi})_{*}\hat{P}\in\Delta(\mathfrak{G}_{\pi})$, is
$\hat{P}_{\mathfrak{G}}(G):=\sum_{u\in(\ell\circ q_{\pi})^{-1}(G)}\hat{P}(u)\qquad(G\in\mathfrak{G}_{\pi}),$
so that
$\hat{P}_{\mathfrak{G}}(G_{\pi,1})=\hat{P}(1)+\hat{P}(2)+\hat{P}(3)=\frac{3}{5},\qquad\hat{P}_{\mathfrak{G}}(G_{\pi,2})=\hat{P}(4)+\hat{P}(5)+\hat{P}(6)=\frac{2}{5}.$
4.2 Kullback–Leibler Divergence
We quantify the information loss caused by coarse-graining via the Kullback–Leibler (KL) divergence between probability distributions due to its well-established role in measuring the divergence between probability distributions in information theory [kullbackleibler1951].
$D_{\text{KL}}(P\parallel Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}$ (1)
While alternative measures such as Jensen-Shannon Divergence or Wasserstein Distance may be explored in future studies, $D_{\text{KL}}$ provides a mathematically straightforward and interpretable baseline for assessing information loss in coarse evaluations.
To illustrate the motivation behind coarse ethics, consider now that we are a small classroom teacher and give a math test to ten students. The results appear in Table 1, Figure 1.
Table 1: Student Scores
Student Score Pass/Fail
$s_{1}$ 5 pass
$s_{2}$ 3 fail
$s_{3}$ 3 fail
$s_{4}$ 3 fail
$s_{5}$ 10 pass
$s_{6}$ 6 pass
$s_{7}$ 2 fail
$s_{8}$ 5 pass
$s_{9}$ 4 fail
$s_{10}$ 0 fail
Figure 1: Distribution of Student Scores (Right-Skewed)
We can represent the results using CGPs as follows: Let the underlying score scale be the finite totally ordered set
$U:=\{0,1,2,\dots,10\}\quad\text{with the natural order }\leq.$
Let the set of students be
$O:=\{s_{1},s_{2},\dots,s_{10}\}.$
Define the score map $f:O\to U$ from Table 1 by
$f(s_{1})=5,\ f(s_{2})=3,\ f(s_{3})=3,\ f(s_{4})=3,\ f(s_{5})=10,$
$f(s_{6})=6,\ f(s_{7})=2,\ f(s_{8})=5,\ f(s_{9})=4,\ f(s_{10})=0.$
The empirical distribution induced by $f$ is $\hat{P}\in\Delta(U)$ defined by
$\hat{P}(u):=\frac{\bigl|\{s\in O:\ f(s)=u\}\bigr|}{|O|}\qquad(u\in U).$
From the score list, the counts are
$\#\{u=0\}=1,\ \#\{u=2\}=1,\ \#\{u=3\}=3,\ \#\{u=4\}=1,$
$\#\{u=5\}=2,\ \#\{u=6\}=1,\ \#\{u=10\}=1,$
and all other scores occur 0 times. Hence
$\hat{P}(0)=\frac{1}{10},\ \hat{P}(1)=0,\ \hat{P}(2)=\frac{1}{10},\ \hat{P}(3)=\frac{3}{10},\ \hat{P}(4)=\frac{1}{10},\$
$\hat{P}(5)=\frac{2}{10},\ \hat{P}(6)=\frac{1}{10},\ \hat{P}(7)=0,\ \hat{P}(8)=0,\ \hat{P}(9)=0,\ \hat{P}(10)=\frac{1}{10}.$
Define a two-grain (interval) partition $\pi$ of $U$ with grains
$G_{\pi,1}:=\{0,1,2,3,4\}\quad(\text{fail-grain}),\qquad G_{\pi,2}:=\{5,6,7,8,9,10\}\quad(\text{pass-grain}),$
and let the associated grain family be
$\mathfrak{G}_{\pi}:=\{G_{\pi,1},G_{\pi,2}\}.$
Let $I_{\pi}:=\{1,2\}$ be the index set and let $L:=\{\text{fail},\text{pass}\}$ be the label set. Define the label map $\ell:I_{\pi}\to L$ by
$\ell(1)=\text{fail},\qquad\ell(2)=\text{pass}.$
Define the coarse assignment map $q_{\pi}:U\to I_{\pi}$ by
$q_{\pi}(u)=\begin{cases}1,&u\in G_{\pi,1},\\ 2,&u\in G_{\pi,2}.\end{cases}$
Then the labeled coarse evaluation is $\ell\circ q_{\pi}:U\to L$, and the induced pass/fail labeling on students is
$\ell\circ q_{\pi}\circ f:O\to L.$
For the given data,
$(\ell\circ q_{\pi}\circ f)(s_{1})=\text{pass},\ (\ell\circ q_{\pi}\circ f)(s_{2})=\text{fail},\ (\ell\circ q_{\pi}\circ f)(s_{3})=\text{fail},\$
$(\ell\circ q_{\pi}\circ f)(s_{4})=\text{fail},\ (\ell\circ q_{\pi}\circ f)(s_{5})=\text{pass},(\ell\circ q_{\pi}\circ f)(s_{6})=\text{pass},\$
$(\ell\circ q_{\pi}\circ f)(s_{7})=\text{fail},\ (\ell\circ q_{\pi}\circ f)(s_{8})=\text{pass},\ (\ell\circ q_{\pi}\circ f)(s_{9})=\text{fail},\$
$(\ell\circ q_{\pi}\circ f)(s_{10})=\text{fail}.$
Define the pushforward distribution on labels by
$\hat{P}_{L}:=(\ell\circ q_{\pi})_{*}\hat{P}\in\Delta(L),\qquad\hat{P}_{L}(\lambda):=\sum_{u\in(\ell\circ q_{\pi})^{-1}(\lambda)}\hat{P}(u)\quad(\lambda\in L).$
Thus,
$\hat{P}_{L}(\text{fail})=\sum_{u=0}^{4}\hat{P}(u)=\frac{6}{10}=\frac{3}{5},\qquad\hat{P}_{L}(\text{pass})=\sum_{u=5}^{10}\hat{P}(u)=\frac{4}{10}=\frac{2}{5}.$
Equivalently, $\hat{P}_{L}(\text{fail})=\hat{P}(G_{\pi,1})$ and $\hat{P}_{L}(\text{pass})=\hat{P}(G_{\pi,2})$.
We compare the probability distributions between the original grading system (ranging from 0 to 10) and the binary pass/fail assessment. More intuitively, if we inform ten students only of the pass/fail criteria and their respective pass or fail status, without disclosing their specific scores—or if we later lose the list of scores and can only recall the pass/fail criteria and each student’s pass or fail status—our task is to determine how much information has been lost.
When using $D_{\text{KL}}$ to achieve this objective, determining $P(i)$ is straightforward. The probability $P(i)$ represents the proportion of students who received a specific score $i$ in the original grading system. It is computed as the ratio of the number of students with score $i$ to the total number of students. It involves assigning the proportion of elements in $O$ that map to each element in $U$. For instance, $P(0)$ is $\frac{1}{10}$, and $P(1)$ is 0 as to Table 1.
4.3 Categorical Unification
The challenge lies in determining $Q(i)$. A straightforward but probably unrealistic assumption would be to assign equal probability to all scores (Table 2). Under the assumption that $Q(i)$ follows a uniform distribution ($Q(i)=\frac{1}{11}$ for all $i$), the $D_{\text{KL}}$ is calculated to be approximately 0.564. This value quantifies the extent to which the uniform assumption deviates from the actual probability distribution $P(i)$.
Table 2: Unification of $Q(i)$
Score Original Probability $P(i)$ Coarse-grained $Q(i)$
0 $\frac{1}{10}$ $\frac{1}{11}$
1 0 $\frac{1}{11}$
2 $\frac{1}{10}$ $\frac{1}{11}$
3 $\frac{3}{10}$ $\frac{1}{11}$
4 $\frac{1}{10}$ $\frac{1}{11}$
5 $\frac{2}{10}$ $\frac{1}{11}$
6 $\frac{1}{10}$ $\frac{1}{11}$
7 0 $\frac{1}{11}$
8 0 $\frac{1}{11}$
9 0 $\frac{1}{11}$
10 $\frac{1}{10}$ $\frac{1}{11}$
However, this assumption is generally unrealistic: even among passing students, scores within the same category (grain) need not be evenly distributed. At the same time, once we coarse-grain the score scale, we want the coarsened distribution to preserve the total probability mass assigned to each grain, and we would like to define a canonical representative distribution that introduces no additional structure beyond this coarse information. To this end, we define Categorical Unification as follows.
Definition 5 (Categorical unification (CU) and $D_{\mathrm{KL\text{-}CU}}$).
Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$. Let $P_{U}$ be a probability distribution on $U$.
(i) The coarse-grained distribution induced by $\pi$ is the pushforward $P_{\pi}$ on $I_{\pi}$ defined by
$P_{\pi}(i):=P_{U}(G_{\pi,i})=\sum_{u\in G_{\pi,i}}P_{U}(u),\qquad i\in I_{\pi}.$
Here, the category index $i$ represents the grain $G_{\pi,i}\subseteq U$; hence $P_{\pi}(i)$ is the total probability mass assigned to that grain, obtained by summing $P_{U}(u)$ over all $u\in G_{\pi,i}$.
(ii) The CU-unification of $P_{U}$ (through $\pi$) is the probability distribution $Q^{\mathrm{CU}}$ on $U$ defined by
$Q^{\mathrm{CU}}(u):=\frac{P_{\pi}(i)}{|G_{\pi,i}|}\qquad\text{for }u\in G_{\pi,i}.$
Here, $Q^{\mathrm{CU}}$ spreads the coarse-grained mass $P_{\pi}(i)$ uniformly across the grain $G_{\pi,i}$; thus each $u\in G_{\pi,i}$ receives probability $Q^{\mathrm{CU}}(u)=P_{\pi}(i)/|G_{\pi,i}|$.
(iii) The KL divergence by categorical unification is
$D_{\mathrm{KL\text{-}CU}}(P_{U}):=D_{\mathrm{KL}}\!\left(P_{U}\,\middle\|\,Q^{\mathrm{CU}}\right),$
where $D_{\mathrm{KL}}(P\|Q)=\sum_{u\in U}P(u)\log\!\frac{P(u)}{Q(u)}$ with the usual convention $0\log(0/q)=0$.
The empirical rationale for this method is as follows: When students are evaluated using a coarse-grained grading system where only categorical results (e.g., pass or fail) are recorded, their exact scores become irretrievable. If a student is classified as “pass”, the only information available is that their score fell within the pass range (e.g., 5 to 10). Without further data, we cannot determine if they barely passed or aced the test. Given this limitation, assigning equal probability to all scores within the pass category is a reasonable and minimally biased assumption. Since no particular score within the pass category has more supporting evidence than another, a uniform assumption ensures neutrality and avoids introducing arbitrary weighting.
The principle of maximum entropy provides a theoretical justification for CU. It states that among all possible probability distributions satisfying a given set of constraints, the one that maximizes entropy (i.e., uncertainty) is the least biased and most informative distribution [jaynes1957]. In the case of pass/fail classification, however, we are not only constrained by the requirement that probabilities sum to one but also by the fact that we know the number of students in each category. Specifically, there are six students in the fail category and four in the pass category, meaning that the total probability for each category is $\frac{6}{10}$ and $\frac{4}{10}$, respectively. According to the principle of maximum entropy, the probability distribution within each category should be uniform to maximize entropy. That is, within the fail category, the assumption that maximizes entropy (and therefore least biased) is a uniform distribution among the five elements (i.e., 0 to 4), and similarly, a uniform distribution should be assumed within the pass category (i.e., 5 to 10). Therefore, CU assigns uniform probability within each category, which corresponds to the entropy-maximizing distribution given the category-level constraints (Table 3). Since an entropy-maximizing distribution is the one that assumes the least additional information, CU can be considered the most unbiased method for probability assignment within coarse-grained categories.
Proposition 2 (Maximum-entropy property of CU).
Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$. Fix $P\in\Delta(U)$ and write
$p_{i}:=P(G_{\pi,i})=\sum_{x\in G_{\pi,i}}P(x)\qquad(i\in I_{\pi}).$
Define the categorical unification (CU) of $P$ (relative to $\pi$) by
$P^{\mathrm{CU}}(x):=\frac{p_{i}}{|G_{\pi,i}|}\quad\text{for }x\in G_{\pi,i}.$
Consider the constraint set
$\mathcal{C}(p,\pi):=\Bigl\{Q\in\Delta(U)\ \bigm|\ Q(G_{\pi,i})=p_{i}\text{ for all }i\in I_{\pi}\Bigr\}.$
Then $P^{\mathrm{CU}}$ maximizes Shannon entropy $H(Q):=-\sum_{x\in U}Q(x)\log Q(x)$ over $\mathcal{C}(p,\pi)$. Moreover, if $p_{i}>0$ for all $i$, the maximizer is unique.
Proof of Proposition 2.
Fix any $Q\in\mathcal{C}(p,\pi)$. For each $i$ with $p_{i}>0$, define the conditional distribution on the grain
$Q_{i}(x):=\frac{Q(x)}{p_{i}}\quad(x\in G_{\pi,i}),$
so that $\sum_{x\in G_{\pi,i}}Q_{i}(x)=1$. A direct calculation yields the entropy decomposition
$H(Q)=H(p)+\sum_{i\in I_{\pi}}p_{i}\,H(Q_{i}),$
where $H(p):=-\sum_{i}p_{i}\log p_{i}$ depends only on $p$ (hence is constant over $\mathcal{C}(p,\pi)$). Therefore, maximizing $H(Q)$ over $\mathcal{C}(p,\pi)$ reduces to maximizing each $H(Q_{i})$ over the simplex $\Delta(G_{\pi,i})$. For a fixed finite set $G_{\pi,i}$, Shannon entropy is maximized uniquely by the uniform distribution on $G_{\pi,i}$, i.e., $Q_{i}(x)=1/|G_{\pi,i}|$ for all $x\in G_{\pi,i}$. Thus, for all $x\in G_{\pi,i}$,
$Q(x)=p_{i}Q_{i}(x)=\frac{p_{i}}{|G_{\pi,i}|}=P^{\mathrm{CU}}(x),$
so $P^{\mathrm{CU}}$ is the (unique, if all $p_{i}>0$) maximizer. ∎
Table 3: Categorical Unification of $Q(i):T=5$
Score Original Probability $P(i)$ Pass/Fail Coarse-grained $Q(i)$
0 $\frac{1}{10}$ Fail $\frac{6}{10}\cdot\frac{1}{5}=\frac{3}{25}$
1 0 Fail $\frac{3}{25}$
2 $\frac{1}{10}$ Fail $\frac{3}{25}$
3 $\frac{3}{10}$ Fail $\frac{3}{25}$
4 $\frac{1}{10}$ Fail $\frac{3}{25}$
5 $\frac{2}{10}$ Pass $\frac{4}{10}\cdot\frac{1}{6}=\frac{1}{15}$
6 $\frac{1}{10}$ Pass $\frac{1}{15}$
7 0 Pass $\frac{1}{15}$
8 0 Pass $\frac{1}{15}$
9 0 Pass $\frac{1}{15}$
10 $\frac{1}{10}$ Pass $\frac{1}{15}$
Theorem 3 (Zero information loss under CU (finite/discrete)).
Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$. Let $P_{U}$ be a probability distribution on $U$, and define the induced coarse-grained distribution
$P_{\pi}(i):=\sum_{u\in G_{\pi,i}}P_{U}(u),\qquad i\in I_{\pi}.$
Define the CU-unification $Q^{\mathrm{CU}}$ on $U$ by
$Q^{\mathrm{CU}}(u):=\frac{P_{\pi}(i)}{|G_{\pi,i}|}\qquad\text{for }u\in G_{\pi,i}.$
Then the discrete KL divergence
$D_{\mathrm{KL}}\!\left(P_{U}\,\middle\|\,Q^{\mathrm{CU}}\right):=\sum_{u\in U}P_{U}(u)\log\frac{P_{U}(u)}{Q^{\mathrm{CU}}(u)}$
is well-defined and satisfies
$D_{\mathrm{KL}}\!\left(P_{U}\,\middle\|\,Q^{\mathrm{CU}}\right)=0\quad\Longleftrightarrow\quad\forall i\in I_{\pi},\ \forall u\in G_{\pi,i}:\ P_{U}(u)=\frac{P_{\pi}(i)}{|G_{\pi,i}|}.$
Equivalently, for every $i\in I_{\pi}$ with $P_{\pi}(i)>0$, the conditional distribution of $P_{U}$ given the event $\{u\in G_{\pi,i}\}$ is uniform on $G_{\pi,i}$:
$P_{U}(\,\cdot\,\mid G_{\pi,i})=\mathrm{Unif}(G_{\pi,i}).$
Proof of Theorem 3.
First note that $D_{\mathrm{KL}}(P_{U}\|Q^{\mathrm{CU}})$ is well-defined. Indeed, if $Q^{\mathrm{CU}}(u)=0$ for some $u\in G_{\pi,i}$, then $P_{\pi}(i)=0$. But $P_{\pi}(i)=\sum_{v\in G_{\pi,i}}P_{U}(v)=0$ implies $P_{U}(v)=0$ for all $v\in G_{\pi,i}$, hence $P_{U}(u)=0$. Therefore $Q^{\mathrm{CU}}(u)=0$ never occurs at a point with $P_{U}(u)>0$.
Now recall the standard fact for the discrete KL divergence: for distributions $P$ and $Q$ on a finite set,
$D_{\mathrm{KL}}(P\|Q)\geq 0,\quad\text{and}\quad D_{\mathrm{KL}}(P\|Q)=0\Longleftrightarrow P=Q.$
Applying this with $P=P_{U}$ and $Q=Q^{\mathrm{CU}}$, we obtain
$D_{\mathrm{KL}}\!\left(P_{U}\,\middle\|\,Q^{\mathrm{CU}}\right)=0\Longleftrightarrow\forall u\in U:\ P_{U}(u)=Q^{\mathrm{CU}}(u).$
By the definition of $Q^{\mathrm{CU}}$, this is equivalent to
$\forall i\in I_{\pi},\ \forall u\in G_{\pi,i}:\ P_{U}(u)=\frac{P_{\pi}(i)}{|G_{\pi,i}|},$
i.e., $P_{U}$ is constant on each grain and equals the grain mass divided equally among its elements. Finally, for any $i$ with $P_{\pi}(i)>0$, the conditional distribution $P_{U}(u\mid G_{\pi,i}):=P_{U}(u)/P_{\pi}(i)$ (for $u\in G_{\pi,i}$) satisfies
$P_{U}(u\mid G_{\pi,i})=\frac{1}{|G_{\pi,i}|},$
which is precisely $\mathrm{Unif}(G_{\pi,i})$. ∎
4.4 Minimizing Information Loss
We aim to make the probability distribution of the coarse-grained pass/fail evaluation as similar as possible to the original probability distribution of scores ranging from 0 to 10. Here, “similar” specifically refers to minimizing information loss, which is quantitatively measured by minimizing the $D_{\text{KL-CU}}$. Suppose we need to divide those ten students (Table 1) into pass and fail groups, but we can determine the passing threshold $(T)$. What score should be used as $T$?
A seemingly natural approach is to use the average score. Since the mean score in this test is 4.1, setting the passing threshold at $T=4$ may seem reasonable. If we classify students scoring below 4 as failing and those scoring 4 or above as passing, we obtain a $D_{\text{KL-CU}}$ of approximately 0.524, which is relatively low. However, this is not the minimum. The lowest $D_{\text{KL-CU}}$ is achieved when the passing threshold is set at $T=7$, resulting in a value of 0.381 (Table 4, Figure 2).
Table 4: $D_{\text{KL-CU}}$ of Table 1
Pass Threshold Pass Students $D_{\text{KL-CU}}$
0 all students 0.564
1 $s_{1},s_{2},s_{3},s_{4},s_{5},s_{6},s_{7},s_{8},s_{9}$ 0.563
2 $s_{1},s_{2},s_{3},s_{4},s_{5},s_{6},s_{7},s_{8},s_{9}$ 0.537
3 $s_{1},s_{2},s_{3},s_{4},s_{5},s_{6},s_{8},s_{9}$ 0.549
4 $s_{1},s_{5},s_{6},s_{8},s_{9}$ 0.524
5 $s_{1},s_{5},s_{6},s_{8}$ 0.521
6 $s_{5},s_{6}$ 0.421
7 $s_{5}$ 0.381
8 $s_{5}$ 0.472
9 $s_{5}$ 0.538
10 $s_{5}$ 0.563
11 no student 0.564
Figure 2: Graphing of Table 4
If an AI teacher performed this calculation and told a human teacher, “Based on this test, only $s_{5}$ should pass”, the human teacher might find it counterintuitive. However, its interpretation is context-dependent. If the number of passing students is fixed (e.g., a predetermined quota of three students), minimizing $D_{\text{KL-CU}}$ may not always be the correct objective. In such cases, the goal is not to preserve the original probability distribution but rather to adhere to external constraints, such as ensuring that exactly three students pass, regardless of their relative performance. This suggests that setting a fixed pass quota can significantly distort the original probability distribution.
Many educators may have encountered situations in which students had widely varying abilities or had to pass students whose performance was lower than they would ideally consider acceptable. This raises a critical question: If the information loss is excessively large, should AI provide an advisory to the teacher? To address this, it may be beneficial to equip the AI teacher with a feature that advises, “The school has mandated that three students must pass this test. Do you believe this number is truly appropriate?” This would allow for human oversight in cases where strict adherence to an imposed quota might compromise fairness or accuracy.
In some contexts, setting the threshold at $T=7$, where only $s_{5}$ passes, may actually align with intuition. Consider a scenario in which the goal is to divide the class into two distinct groups: one for mathematically advanced students (e.g., an honors class) and another for those who need more support. The original score distribution implies that only one student significantly outperforms the others in mathematics; accordingly, $T=7$ identifies a single clear outlier ($s_{5}$). Thus, setting $T=7$ produces a classification that best preserves the structure of the original probability distribution while maintaining a meaningful distinction between high and low performers.
This analysis highlights how the calculation of $D_{\text{KL-CU}}$ can serve as a bridge between AI-driven decision-making and human intuition, offering a foundational example for both XAI and CE.
5 Discussion
5.1 Mathematical Foundation of Coarse Ethics
CGPs have demonstrated that the rationality of coarse evaluations, as argued by CE, can be rigorously analyzed. The coarse evaluations we use in daily life—for instance, the assessment that “Olivia received an excellent grade on her mathematics test”—exhibit a structured mathematical form within CGPs. The fact that an evaluation is coarse—for example, that Olivia’s specific score difference from other students who also received an “excellent” rating is intentionally disregarded—does not imply that the evaluation is ambiguous. Here, we can clearly distinguish between coarseness and ambiguity in assessment.
Traditional ethical theories, particularly consequentialist (e.g., felicific calculus by Bentham [bentham1789introduction]) and virtue-based approaches (e.g., Stoicism), often emphasize continuity in moral evaluation, where actions that align more closely with a given ideal are ranked proportionally higher. However, CE suggests that even when individuals differ significantly in their alignment with an ideal, they may still be evaluated equivalently [izumoweng2022]. This aligns with Immanuel Kant’s criticism of excessive ethical precision, particularly his rejection of casuistry—the attempt to make moral laws overly specific through case-by-case analysis [kant2017]. A simple example illustrates this point: Suppose a student receives a failing grade of 30 points on an essay from a human teacher. Dissatisfied, they submit the same essay to an AI teacher for evaluation, which assigns them a score of 30.1 while their peer, who was also given 30 by the human teacher, receives 29.8 from the AI. If the student then argues that the AI’s refined scoring justifies a reconsideration of their grade, such a claim would be unreasonable. From the teacher’s perspective, both evaluations reflect insufficient performance, and fixating on a marginal difference of 0.3 points does not meaningfully change the outcome.
5.2 Tradeoff Between Evaluation Efficiency and Information Retention
As demonstrated in this study, when $Q(i)$ in $D_{\text{KL}}$ is processed through categorical unification, information is preserved only in specific cases. Our findings suggest a potential trade-off between evaluation efficiency and information retention. While this study does not provide a formal proof, we hypothesize that as evaluation processes prioritize simplicity and accessibility, the loss of fine-grained details becomes inevitable. Further research could investigate whether there exists a Pareto-optimal balance between evaluation efficiency and retained information in AI-driven decision-making.
Assuming a uniform distribution within a category is the most reasonable default when no additional information is available. When only a pass label is recorded, any distinction between different passing scores is inherently erased, making a categorically uniform distribution the simplest assumption that reflects this lack of distinguishing data. While alternative assumptions, such as normal or skewed distributions, might seem more refined, they require external data that is often unavailable in real-world grading systems. If historical trends suggest a particular distribution (e.g., most passing students score near the middle of the range), Bayesian updates could provide a better-informed probability model. However, in the absence of such priors, assuming a categorically uniform distribution remains the least biased approach.
5.3 Application to Explainable AI
We apply the above discussion to the challenges identified in Section 1.2. One key application is in refining AI-driven safety advisories. For example, in autonomous driving systems, the coarse-graining of safety warnings must not result in the loss of critical risk information. Instead of issuing a generic warning such as “Your driving is dangerous”, CGPs can guide AI systems to adaptively determine the level of warning granularity, ensuring that important details about immediate risks (e.g., speed excess relative to road conditions or proximity to nearby obstacles) are retained within the coarsened message. This approach allows for a structured balance between reducing cognitive overload and preserving necessary safety information.
Additionally, CGPs offer a refined approach to addressing the limitations of the three-tiered explanatory model in CE. Rather than strictly categorizing explanations into those for experts, general users, and individuals with cognitive difficulties, CGPs allow for dynamic adaptation of explanation granularity by integrating user profiling mechanisms. In a medical diagnostic AI, for example, an expert might receive a detailed probability breakdown of differential diagnoses, while a general patient would be presented with a high-level summary emphasizing key risks and actionable next steps. This ensures that explanations remain accessible without sacrificing necessary details for informed decision-making. This personalized granularity adjustment can improve the effectiveness of AI explanations in domains such as medical diagnostics, legal decision-making, and financial advising, where different users require distinct levels of detail in their assessments.
6 Conclusion
This study has established a rigorous mathematical foundation for Coarse Ethics (CE). By introducing the framework of CGPs, we demonstrated that coarse evaluations can be systematically analyzed rather than being treated as vague or arbitrary assessments. A coarse set is defined by constructing a coarse-grained partition over a totally ordered set, and its mathematical properties have been formalized through the definitions, propositions, and theorems in this paper. The key conclusions drawn from this study are as follows:
  1. 1.
    Coarse evaluation methods, as described in CE, can be formulated with mathematical precision, providing a structured approach to assessments.
  2. 2.
    Different coarse evaluation methods can be analyzed within the framework of CGPs. In particular, each evaluation method can be represented as a probability distribution, allowing for probabilistic and statistical comparison.
  3. 3.
    A quantitative approach to evaluating information loss in coarse assessments is provided by Kullback-Leibler Divergence, which measures the deviation of a coarse-grained distribution from the original evaluation.
As AI systems become more complex, the need for interpretable and adaptable explanations grows. Current XAI techniques, such as LIME and SHAP, often provide explanations at a fixed granularity, which may not align with user-specific cognitive requirements. CGPs offer a structured way to dynamically adjust the level of detail in AI explanations, ensuring that different users receive information at an appropriate level of abstraction. This approach could enhance AI-assisted decision-making in critical fields such as medicine, finance, and law, where balancing interpretability and complexity is essential.
This study presents an initial exploration of CGPs, but several limitations remain. Our analysis is restricted to totally ordered sets and assumes a covering condition, meaning that every element of the original set is included within the coarse-grained structure. This differs from previous research, which allows for the relaxation of coverage constraints. In other words, our approach focuses on cases where a coarse evaluation fully encompasses the set of evaluated entities. Additionally, our discussion has been primarily framed within the context of educational assessments, particularly school testing. Future research should investigate the implications of relaxing the covering condition and explore applications in broader ethical and legal decision-making frameworks.
Appendix A. Python Code for Calculating $D_{\text{KL-CU}}$
Listing 1: Python Code for $D_{\text{KL-CU}}$ of Table 4
import numpy as np
import matplotlib.pyplot as plt
from collections import Counter
def compute_probability_distribution(scores):
"""ComputeP(i)=freq(i)/Nforscores0..max(scores)."""
total = len(scores)
max_score = max(scores)
# Include missing scores with 0 frequency
freq = Counter(scores)
p = {i: freq.get(i, 0) / total for i in range(max_score + 1)}
return p
def coarse_grain_distribution(p, threshold):
"""GrouppbyTanddistributeprobabilityuniformly."""
scores = sorted(p.keys())
fail_group = [s for s in scores if s < threshold]
pass_group = [s for s in scores if s >= threshold]
q = {}
if fail_group:
total_fail = sum(p[s] for s in fail_group)
for s in fail_group:
q[s] = total_fail / len(fail_group)
if pass_group:
total_pass = sum(p[s] for s in pass_group)
for s in pass_group:
q[s] = total_pass / len(pass_group)
return q
def kl_divergence(p, q):
"""ComputeD_KL(P||Q)=sumP(i)*log(P(i)/Q(i))."""
dkl = 0
for s in p:
if p[s] > 0 and q.get(s, 0) > 0:
dkl += p[s] * np.log(p[s] / q[s])
return dkl
def analyze_thresholds(scores):
"""ComputePandforTcomputeQandD_KL.
␣␣␣␣Returnslistof(T,D_KL)."""
p = compute_probability_distribution(scores)
max_score = max(scores)
results = []
# Iterate T from 0 to max_score+1
for T in range(max_score + 2):
q = coarse_grain_distribution(p, T)
dkl = kl_divergence(p, q)
results.append((T, dkl))
return results
# Test data
test_scores = [5, 3, 3, 3, 10, 6, 2, 5, 4, 0]
# Compute and print KL divergence for each threshold T
results = analyze_thresholds(test_scores)
for T, dkl in results:
print(f"T={T:2d}␣␣–>␣␣D_KL(P||Q)={dkl:.5f}")
# Find optimal threshold with minimal KL divergence
best_threshold, min_dkl = min(results, key=lambda x: x[1])
print(
[…]
f"minimumKLdivergenceisapprox.{min_dkl:.5f}"
# Plot the results (optional)
Ts, dkl_vals = zip(*results)
plt.figure(figsize=(8, 5))
plt.plot(Ts, dkl_vals, marker=’o’, linestyle=’-’)
plt.xlabel(’Pass/FailThreshold(T)’)
plt.ylabel(’KLDivergenceD_KL(P||Q)’)
plt.title(’KLDivergenceforEachThreshold’)
plt.grid(True)
plt.show()
Declarations
Funding
This research received no specific grant from any funding agency in the public, commercial, or not-for-profit sectors.
Conflicts of Interest
The author declares no conflict of interest.
Ethical Approval
This article does not contain any studies with human participants or animals performed by any of the authors.
Data Availability
No empirical data was used in this study.
Code Availability
The Python code used to compute the KL Divergence is available at https://github.com/Takashi-Izumo/CGPs-implementation.

### v7

@@
-Information Loss under Coarse-Grained Partitions: A Discrete Framework for AI Ethics
+Quantifying Information Loss under Coarse-Grained Partitions: A Discrete Framework for Explainable Artificial Intelligence
 Abstract
-As artificial intelligence (AI) systems become increasingly embedded in ethically sensitive domains such as education, healthcare, and transportation, the need to balance accuracy and interpretability in decision-making has become a central concern.
-Coarse Ethics (CE) motivates the use of coarse-grained evaluations, such as letter grades, as ethically appropriate under cognitive, institutional, and contextual constraints.
-However, CE has lacked a simple mathematical formalization that makes the informational cost of such coarse judgments explicit.
-This paper introduces coarse-grained partitions (CGPs) as a discrete, information-theoretic framework for modeling coarse evaluation.
-We treat an underlying score scale as a finite ordered set, represent coarse evaluation as a partition into grains together with an index assignment, and define the induced coarse-grained distribution by pushforward.
-To quantify the informational impact of coarse-graining, we propose categorical unification (CU), which constructs a canonical within-grain unification of the original distribution, and we measure information loss by the discrete Kullback–Leibler divergence between the original distribution and its CU-unification (KL-CU).
-We show that information loss is zero if and only if the original distribution is already uniform within each grain, clarifying precisely when coarse evaluation preserves all relevant information.
-We illustrate the framework with educational grading and explainable AI (XAI) settings, where coarse categories are often required for communication and accountability.
-By providing a tractable discrete formalization of coarse evaluation and a principled measure of information loss, CGPs offer a foundation for designing interpretable AI systems that balance comprehensibility, fairness, and informational integrity.
+As artificial intelligence (AI) systems are increasingly used in ethically sensitive domains such as education, healthcare, and transportation, balancing accuracy and interpretability has become a central concern.
+Coarse ethics (CE) motivates coarse-grained evaluations under cognitive, institutional, and contextual constraints, but it still lacks a simple mathematical formalization of admissible coarse-graining and its informational consequences.
+This paper introduces coarse-grained partitions (CGPs) as a discrete framework for modeling coarse evaluation on a finite totally ordered score scale.
+A CGP represents coarse evaluation as a partition into grains with an index assignment, and induces a coarse-grained distribution by pushforward.
+To compare admissible coarse-grainings, we introduce categorical unification (CU), which constructs a canonical fine-scale reconstruction from the coarse representation under minimal assumptions.
+On this basis, we define a KL-based measure of information loss, $D_{\mathrm{KL\text{-}CU}}$, as the divergence between the original fine-grained distribution and its CU-based reconstruction.
+We prove that $D_{\mathrm{KL\text{-}CU}}=0$ if and only if the original distribution is already uniform within each grain.
+This shows that zero loss, in the sense of the proposed measure, is a highly exceptional limiting case rather than a realistic benchmark for ordinary evaluative practice.
+We also show that the framework leads naturally to an optimization problem for comparing alternative admissible CGPs.
+Applications to educational grading and explainable AI (XAI) illustrate how the framework clarifies trade-offs among informational fidelity, interpretability, and coarsening cost.
 keywords:
@@
 Techniques such as Local Interpretable Model-Agnostic Explanations (LIME, [ribeiro2016whyitrustyou]) and SHapley Additive exPlanations (SHAP, [lundberg2017unifiedapproachinterpretingmodel]) have been widely used to provide post-hoc interpretability for AI systems.
-From an ethical standpoint, Coarse Ethics (CE, [izumoweng2022]) proposes that coarse-grained assessments are not only necessary but also ethically justifiable in many decision-making contexts.
-For example, academic grading systems classify students into letter grades or grade point average (GPA), abstracting away precise score differences.
-Similarly, phrases like “Olivia and Noah are Harvard graduates” generalize individual performance details.
+Recent research on AI-assisted decision-making has increasingly emphasized that explanation and evaluation are not purely technical matters, but also ethical ones.
+In particular, a growing body of work has examined how AI outputs should be presented to human users in ways that are cognitively manageable, socially appropriate, and normatively defensible.
+Several prior studies are relevant in this respect.
+For example, He et al.
+[he2024analogies] explored how analogies can assist laypeople in AI-assisted decision-making, highlighting the cognitive conditions under which explanations become understandable to non-expert users.
+Peters and Carman [peters2024cultural] examined the influence of cultural bias in XAI research, showing that users from different cultural backgrounds may interpret AI explanations differently.
+These studies underscore that the form and granularity of AI explanations are not neutral, but depend on the capacities, expectations, and interpretive contexts of their audiences.
+From this broader ethical perspective, coarse ethics (CE, [izumoweng2022], [izumo2025]) offers a particularly suggestive framework.
+CE proposes that coarse-grained assessments are not only practically unavoidable but also ethically justifiable in many decision-making contexts.
+For example, academic grading systems classify students into letter grades or grade point averages (GPAs), thereby abstracting away precise score differences, and expressions such as “Olivia and Noah are Harvard graduates” similarly suppress finer distinctions in individual performance.
 In AI-driven decision-making, CE argues that the legitimacy of transforming fine-grained evaluations into coarse ones depends on the cognitive capacity and contextual needs of the audience [izumo2025].
-1.2 Challenges in Coarse Ethics
-Despite its conceptual appeal, CE faces two primary challenges:
-  1. 1.
-    Lack of Mathematical Rigor: While CE provides an ethical justification for coarse evaluations, it lacks a mathematically formalized framework.
-For instance, Izumo [izumo2025] proposes that an AI assistant should issue coarse warnings such as “Your driving is dangerous” rather than continuously recalculating and displaying the optimal speed.
-However, the study does not specify the exact computational method to determine when and how such warnings should be issued.
-Without a formalized structure, these decisions may be inconsistent, raising concerns about reliability in AI-driven risk assessment.
-  2. 2.
-    Ambiguity in Categorization: Prior research has introduced a three-tiered explanatory approach—detailed explanations for experts, coarse explanations for general users, and even coarser explanations for individuals with cognitive limitations [izumo2025].
-However, this classification is overly simplistic and lacks adaptability, failing to account for the diverse range of cognitive and contextual factors in human decision-making.
-Addressing these limitations is essential, particularly as AI advances toward artificial general intelligence (AGI).
-Strengthening the mathematical foundation of CE and refining its categorization mechanisms can contribute to the development of more robust, context-aware AI explanations.
+The present paper focuses on this CE perspective.
+Despite its conceptual appeal, CE faces a central challenge: it lacks a mathematically rigorous structure, and this weakens the usefulness of the very notion of coarse-grained normative evaluation.
+Izumo and Weng [izumoweng2022] identify two requirements for coarsening a fine-grained evaluation: first, that the coarse-grained evaluation should sufficiently cover the objects evaluated in the original one (for example, if Olivia is evaluated in the original assessment, she should not disappear from the coarse-grained version); and second, that the coarse-grained evaluation should not reverse the value ordering embodied in the fine-grained one (for example, if Olivia is rated higher than Noah in the original evaluation, the coarsened evaluation should not end up ranking Noah above Olivia).
+However, as the authors themselves acknowledge, these conditions do not suffice to determine a unique coarse-grained evaluation.
+As a result, multiple distinct coarse evaluations satisfying those conditions may arise from a single fine-grained one.
+This problem stems from the fact that the previous literature does not specify the mathematical process by which coarsening is carried out.
 1.2 Methodology
-Several prior studies share conceptual similarities with CE.
-For example, He et al.
-[he2024analogies] explored how analogies can assist laypeople in AI-assisted decision-making.
-Their work emphasizes the cognitive aspects of explanation reception and user comprehension.
-While our study shares the objective of improving human-AI interaction, we focus on the mathematical quantification of information loss in coarse evaluations rather than the psychological mechanisms of explainability.
-Peters and Carman [peters2024cultural] examined the influence of cultural bias in XAI research, demonstrating how users from different cultural backgrounds interpret AI explanations differently.
-Their study focuses on the subjective perception of AI explanations rather than developing a mathematical framework for evaluation.
-In contrast, our work formalizes coarse-grained evaluations using set theory, providing a structured approach to managing granularity in AI-driven decision-making.
-This paper proposes coarse-grained partitions (CGPs) as a novel set-theoretic approach to resolving these issues.
-Set theory has been extensively applied to human decision-making, forming the basis of fuzzy set theory [zadeh1965fuzzy], rough set theory [pawlak1982rough], and soft set theory [molodtsov1999soft].
-These approaches primarily focus on modeling real-world uncertainty through mathematical representations.
-In contrast, CGPs shift the focus from enhancing AI’s ability to analyze the world to helping humans extract and interpret AI-generated insights.
-While previous works aimed at improving AI’s analytical capabilities with set theory [10.5120/8924-2996, 9015977, dalkilic2022novel], CGPs are based on the premise that AI models now outperform human experts in specific domains, as demonstrated by recent studies on AI surpassing human-level performance in standardized exams [bicknell2024chatgpt, katz2023gpt4].
-Consequently, the key question is no longer how AI perceives reality, but rather how humans can efficiently understand and interact with AI-driven decisions.
-This perspective is particularly relevant as Large Language Models (LLMs) and domain-specific AI systems continue to exceed human performance in complex tasks [vaswani2017attention, radford2018improving, brown2020language, chowdhery2022palm].
-As AI models continue to outperform human experts in specialized tasks, discussions surrounding AGI expectations and ethical considerations have intensified [morris2024levelsagioperationalizingprogress].
-By establishing a formal mathematical foundation for coarse evaluations, CGPs offer a critical framework for addressing these challenges and redefining the evolving relationship between AI and human cognition.
-2 Coarse-Grained Partitions
+This paper proposes coarse-grained partitions (CGPs) as a novel set-theoretic framework for addressing these issues.
+Set-theoretic approaches have long been used in the study of human decision-making, most notably in fuzzy set theory [zadeh1965fuzzy], rough set theory [pawlak1982rough], and soft set theory [molodtsov1999soft].
+These approaches were primarily developed to represent uncertainty, vagueness, and incomplete information in the world as mathematically tractable structures.
+CGPs differ from these approaches in their point of departure.
+Rather than improving AI’s ability to model external reality, CGPs are designed to analyze how fine-grained evaluations can be transformed into coarser forms that remain usable for human agents.
+In this sense, the problem is no longer how AI perceives the world, but how human beings can understand, interpret, and act upon AI-generated evaluations.
+This perspective becomes especially important in domains where AI systems now outperform human experts in narrowly defined tasks, including standardized testing and other forms of specialized assessment [bicknell2024chatgpt, katz2023gpt4].
+As AI systems become increasingly capable, the challenge of translating their internal evaluations into forms intelligible to human users becomes correspondingly more urgent.
+From this perspective, CGPs provide a mathematical framework for studying coarse-grained evaluation itself.
+We formalize the transformation from fine-grained to coarse-grained evaluation by introducing partitions of a finite discrete totally ordered score scale.
+We then compare the probability distributions before and after coarsening, define a canonical reconstruction of the fine-grained distribution through categorical unification, and measure information loss by KL divergence [kullbackleibler1951] with categorical unification (CU), i.e., $D_{\mathrm{{KL\text{-}CU}}}$.
+This allows us to analyze, in precise terms, when coarse-graining preserves information, when it necessarily destroys it, and how alternative coarse-graining schemes may be compared.
+More precisely, the role of CGPs in this paper is twofold.
+First, as interval partitions of a finite totally ordered scale, they provide a mathematically explicit class of coarse-grainings that satisfies the structural requirements emphasized in CE, namely coverage and order preservation.
+Second, since these requirements alone do not determine a unique coarse-graining, we introduce $D_{\mathrm{{KL\text{-}CU}}}$ not as a replacement for them, but as an additional criterion for comparing admissible CGPs in terms of informational loss.
+The structure of the paper is as follows.
+Section 2 introduces the basic framework of coarse-grained partitions on finite discrete totally ordered sets.
+Section 3 formalizes the object-to-category map and the induced probability distributions associated with coarse-grained evaluation.
+Section 4 defines CU and the KL-based measure of information loss, and proves the zero-information-loss theorem.
+Section 5 examines the optimization problem of coarse-graining design and discusses its significance for explainable AI and related decision-making contexts.
+Section 6 concludes.
+2 Coarse-Grained Evaluation
+2.1 Coarse-Grained Partitions
+In many institutional settings, normative evaluations are implemented on finite, discrete score sets.
+For instance, school test scores are recorded as bounded integers rather than real numbers.
+In this paper, we take such discretized scores as our main motivating case.
+Accordingly, it suffices to model the fine-grained score domain as an arbitrary finite totally ordered set (e.g., a finite subset of $\mathbb{N}$).
 Throughout the paper, we fix a finite totally ordered set $(U,\leq)$, which we call the underlying scale.
@@
 We equip $U$ with the discrete $\sigma$-algebra $2^{U}$.
-2.1 Coarse-Grained Partitions
+What must be discussed rigorously, when we speak of coarsening a fine-grained evaluation, is how the resulting coarse-graining should be represented as a mathematical object.
+Here we formalize it as a coarse-grained partition of $U$, which will later induce a score-to-category map.
 Definition 1 (Coarse-grained partition).
-Let $(U,\leq)$ be a finite totally ordered set, and let $\mathscr{I}(U,\leq)$ denote the family of all nonempty order-convex (interval) subsets of $U$.
+Given the underlying scale $(U,\leq)$, let $\mathscr{I}(U,\leq)$ denote the family of all nonempty intervals of $U$.
 A coarse-grained partition of $(U,\leq)$ is a partition object $\pi$ equipped with a finite index set $I_{\pi}$ and a family of grains
@@
 We call each $G_{\pi,i}$ a grain (or category) of $\pi$.
-Here $\pi$ denotes the partition object, while $\mathfrak{G}_{\pi}$ denotes its underlying family of grains.
-Remark 1 (On generality).
-The discrete ordered framework above is sufficient for the main results of this paper and matches typical applications such as grading scales.
-A measure-theoretic extension to general measurable spaces can be developed separately (e.g.
-by replacing sums with integrals), and may be placed in an appendix if needed.
-Remark 2 (Ordering grain families).
-A key structural feature of interval coarse-graining is that it induces an order not only on $U$ but also on the grain family itself. Define, for $G,H\in\mathfrak{G}_{\pi}$,
-$G\preccurlyeq H\quad:\Longleftrightarrow\quad(\forall x\in G)(\forall y\in H)\;x\leq y.$
-This element-wise relation captures a global ordering of categories and is compatible with the intuitive notion that some coarse-grained categories are uniformly “lower” (or “worse”) than others along the evaluation scale.
-Although the present paper focuses on probabilistic pushforwards and information loss under within-grain unification, this induced grain order provides the missing structure needed to define order-sensitive operators on partitions (e.g., adjacent merging, monotone coarsening, and order-constrained optimization of coarse-grainings).
-Remark 3 (No internal order within grains).
-In our constructions, each grain $G\in\mathfrak{G}_{\pi}$ is treated as an unordered set: only membership $u\in G$ and the cardinality $|G|$ are used.
-In particular, the CU construction assigns probabilities uniformly within each grain, so no internal ordering of elements inside a grain is required in this paper.
-By contrast, in extensions where one selects a representative value for each grain—for instance, taking $\min(G)$ as the representative—the ambient order on $U$ (and hence the induced order within $G$) becomes relevant.
+Here $\pi$ denotes the structure specifying how $U$ is partitioned, while $\mathfrak{G}_{\pi}$ denotes its family of grains.
 Example 1.
@@
 In this context, each element of the grain family associated with a partition (e.g., $\{1,2\}$ in $\mathfrak{G}_{\pi_{2}}$) is a grain (interval) of $U_{1}$, so this terminology is natural.
-The collection of (interval) coarse-grained partitions of $U$ is denoted by $\mathcal{P}_{\textrm{int}}(U)$ in this paper. In particular, for the example $U_{1}=\{1,2,3\}$ above, we have
-$\mathcal{P}_{\textrm{int}}(U_{1})=\{\pi_{1},\pi_{2},\pi_{3},\pi_{4}\},$
-where each partition $\pi_{k}$ is represented by its grain family $\mathfrak{G}_{\pi_{k}}$ as listed in the example. Accordingly, the collection of grain families induced by interval partitions is
-$\mathscr{G}_{\textrm{int}}(U_{1})=\{\mathfrak{G}_{\pi_{1}},\mathfrak{G}_{\pi_{2}},\mathfrak{G}_{\pi_{3}},\mathfrak{G}_{\pi_{4}}\}.$
+Remark 1 (Ordering grain families).
+A key structural feature of interval coarse-graining is that it induces an order not only on $U$ but also on the grain family itself. Define, for $G,H\in\mathfrak{G}_{\pi}$,
+$G\preccurlyeq H\quad:\Longleftrightarrow\quad(\forall x\in G)(\forall y\in H)\;x\leq y.$
+This element-wise relation captures a global ordering of categories and is compatible with the intuitive notion that some coarse-grained categories are uniformly “lower” (or “worse”) than others along the evaluation scale.
+Although the present paper focuses on probabilistic pushforwards and information loss under within-grain unification, this induced grain order provides the missing structure needed to define order-sensitive operators on partitions (e.g., adjacent merging, monotone coarsening, and order-constrained optimization of coarse-grainings).
 If arbitrary set partitions were allowed, then the number of partitions of an $n$-element set would be the Bell number $B_{n}$, which grows extremely rapidly.
@@
 ∎
-3 Coarse Assignment and Category Assignment
-3.1 Coarse Assignment
+3 Object-to-Category Map
+3.1 Definition
 CE studies how a fine-grained evaluation on a scale $U$ is mapped into a coarser categorization.
+In practice, an evaluation is always performed of some collection of objects.
+For example, the evaluation “Emily is in rank A because she scored 90 on a math test” can be viewed as a two-step mapping: Emily (as an object) is first mapped to a fine-grained numerical score (90 points), and this score is then mapped to a coarse category (rank A).
+Accordingly, we need a map from each object to a category (an object-to-category map); however, this map is the composition of two maps: one from the object to a fine-grained score (an object-to-score map), and another from that fine-grained score to a coarse category (a score-to-category map).
+We begin by defining the score-to-category map.
 Given a coarse-grained partition $\pi$ of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$, the coarse categories are conceptually given by the grains $G_{\pi,i}\in\mathfrak{G}_{\pi}$.
-For technical convenience (measurability and pushforward constructions), we represent these categories by their labels $I_{\pi}$ and fix a coarse assignment map
+For technical convenience (measurability and pushforward constructions), we represent these categories by their labels $I_{\pi}$ and fix a score-to-category map
 $q_{\pi}:U\to I_{\pi},\qquad q_{\pi}(x)=i\ \Longleftrightarrow\ x\in G_{\pi,i}.$
-Definition 2 (Index set and coarse assignment map).
+Definition 2 (Score-to-category map).
 Let $(U,\leq)$ be a finite totally ordered set, equipped with the discrete $\sigma$-algebra $2^{U}$.
-Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ (cf.
+Let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ (cf.
 Def.
 1).
-Define the index set by
-$I_{\pi},$
-equipped with the discrete $\sigma$-algebra $2^{I_{\pi}}$. The coarse assignment map associated with $\pi$ is the map $q_{\pi}:U\to I_{\pi}$ defined by
+Define the index set $I_{\pi}$ as the label set appearing in $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$, equipped with the discrete $\sigma$-algebra $2^{I_{\pi}}$.
+The score-to-category map associated with $\pi$ is the map $q_{\pi}:U\to I_{\pi}$ defined by
 $q_{\pi}(x)=i\quad\Longleftrightarrow\quad x\in G_{\pi,i}.$
-Example 2 (A coarse assignment map).
+Example 2 (A score-to-category map).
 Let $U=\{1,2,3\}$ with the discrete $\sigma$-algebra $2^{U}$. Define a coarse-grained partition $\pi$ of $U$ by the grain family
 $\mathfrak{G}_{\pi}:=\{G_{\pi,i}\}_{i\in I_{\pi}}=\bigl\{\{1\},\{2,3\}\bigr\},\qquad I_{\pi}:=\{1,2\}.$
-Define the coarse assignment map $q_{\pi}:U\to I_{\pi}$ by $q_{\pi}(u)=i$ iff $u\in G_{\pi,i}$. Then
+Define the score-to-category map $q_{\pi}:U\to I_{\pi}$ by $q_{\pi}(u)=i$ iff $u\in G_{\pi,i}$. Then
 $q_{\pi}(1)=1,\qquad q_{\pi}(2)=2,\qquad q_{\pi}(3)=2.$
-Remark 4 (Measurability in the finite/discrete setting).
-Since we work with the discrete $\sigma$-algebras $2^{U}$ and $2^{I_{\pi}}$ on finite sets, every map between these spaces is measurable.
-In particular, the coarse assignment map $q_{\pi}:U\to I_{\pi}$ is automatically measurable.
-3.2 Object-to-Category Map
-In practice, an evaluation is always performed of some collection of objects.
-Therefore, it is not enough to specify only the relationship between a fine-grained scale and its coarse-grained counterpart.
-One must also fix (i) how objects are mapped to fine-grained scores, and (ii) how those scores are subsequently mapped to coarse categories.
-In other words, we must specify both the object-to-score map and the induced object-to-category assignment.
 Definition 3 (Object-to-category map).
 Let $O$ be a nonempty set of objects and let $f:O\to U$ be an object-to-score map.
-Fix a coarse-grained partition $\pi$ of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ and coarse assignment map $q_{\pi}:U\to I_{\pi}$.
+Fix a coarse-grained partition $\pi$ of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ and score-to-category map $q_{\pi}:U\to I_{\pi}$.
 The object-to-category map induced by $f$ and $\pi$ is the composition
@@
 4.1 Probability Measures
-Since normative assessments are typically made over a finite set of cases in practice, this section restricts attention to the finite setting.
-We model both objects and scores as finite sets and represent empirical frequencies by discrete probability distributions.
-This choice keeps the information-loss analysis entirely within the standard discrete (Shannon/KL) formalism, without comparing probability measures of different types (e.g., continuous vs.
-atomic).
+Since the previous section defined the evaluation objects and the maps $f:O\to U$ (from objects to fine-grained scores) and $q_{\pi}:U\to I_{\pi}$ (from fine-grained scores to category labels), we now compare the induced distributions on the fine-grained score scale $U$ and on the label set $I_{\pi}$.
+Concretely, given a probability distribution on $O$, we obtain the score distribution on $U$ as the pushforward along $f$, and the coarse distribution on $I_{\pi}$ as the pushforward along $q_{\pi}\circ f$.
 Definition 4 (Pushforward (coarse-grained) probability).
-Let $U$ be a finite set and equip it with the power set $\Sigma:=2^{U}$.
-Let $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$ be a finite coarse-grained partition of $U$, and let $P_{U}$ be a probability distribution on $(U,2^{U})$.
-With $q_{\pi}$ as in Definition 2, define the coarse-grained probability distribution $P_{\pi}$ on $(I_{\pi},2^{I_{\pi}})$ by the pushforward
-$P_{\pi}:=P_{U}\circ q_{\pi}^{-1}.$
-Equivalently, for each $i\in I_{\pi}$,
-$P_{\pi}(\{i\})=P_{U}(G_{\pi,i}).$
-Remark 6 (The coarse $\sigma$-algebra generated by $\pi$).
-The partition $\pi$ induces the $\sigma$-algebra
-$\Sigma_{\pi}:=\sigma(\pi)=\Bigl\{\bigcup_{i\in J}G_{\pi,i}\;:\;J\subseteq I_{\pi}\Bigr\}\subseteq\Sigma.$
-Moreover, $q_{\pi}$ is $(\Sigma,\;2^{I_{\pi}})$-measurable and $\Sigma_{\pi}=q_{\pi}^{-1}(2^{I_{\pi}})$.
-The space $O$ represents the underlying domain of objects (e.g., test scores), equipped with a probability distribution $\mathbb{P}$ that captures how likely different objects are to occur.
-The map $f:O\to U$ assigns each object in $O$ to a category in $U$.
-Definition 4 defines the induced probability distribution on $U$ by pushing forward $\mathbb{P}$ through $f$: for any measurable set of categories $A\subseteq U$, the probability that the assigned category lies in $A$ is defined as the probability mass of all objects in $O$ that are mapped into $A$, i.e.,
-$\mathbb{P}_{U}(A)=\mathbb{P}(f^{-1}(A)).$
-In other words, $\mathbb{P}_{U}$ is the distribution of the categorized outcome $f(o)$ when $o$ is drawn according to $\mathbb{P}$.
-Example 3 (Pushforward to a win/lose outcome).
-As a running toy example, consider a dice game in which a larger face value yields a higher score. We take the underlying scale to be
-$U:=\{1,2,3,4,5,6\}\quad\text{equipped with the natural total order }1\leq 2\leq\cdots\leq 6.$
-Let $O:=\{a,b,c,d,e\}$ be a set of five players (or five trials), and define an outcome map $f:O\to U$ by
-$f(a)=1,\quad f(b)=1,\quad f(c)=3,\quad f(d)=5,\quad f(e)=6.$
-The induced empirical distribution $\hat{P}\in\Delta(U)$ is defined by
-$\hat{P}(u):=\frac{|\{o\in O:f(o)=u\}|}{|O|}\qquad(u\in U).$
-Hence,
-$\hat{P}(1)=\frac{2}{5},\ \hat{P}(2)=0,\ \hat{P}(3)=\frac{1}{5},\ \hat{P}(4)=0,\ \hat{P}(5)=\frac{1}{5},\ \hat{P}(6)=\frac{1}{5}.$
-Next, define a two-grain partition $\pi$ of $U=\{1,2,3,4,5,6\}$ by the grain family
-$G_{\pi,1}:=\{1,2,3\},\qquad G_{\pi,2}:=\{4,5,6\},$
-and denote the associated grain family by
-$\mathfrak{G}_{\pi}:=\{G_{\pi,1},G_{\pi,2}\}.$
-(For readability, we may refer to $G_{\pi,1}$ as the Lose-grain and $G_{\pi,2}$ as the Win-grain.)
-Let $I_{\pi}:=\{1,2\}$ and define the coarse assignment map $q_{\pi}:U\to I_{\pi}$ by
-$q_{\pi}(u)=\begin{cases}1,&u\in G_{\pi,1}=\{1,2,3\},\\ 2,&u\in G_{\pi,2}=\{4,5,6\}.\end{cases}$
-Introduce the grain-label map $\ell:I_{\pi}\to\mathfrak{G}_{\pi}$ by
-$\ell(1)=G_{\pi,1},\qquad\ell(2)=G_{\pi,2}.$
-Then $\ell\circ q_{\pi}:U\to\mathfrak{G}_{\pi}$ maps each outcome $u\in U$ to its grain, and hence
-$\ell\circ q_{\pi}\circ f:O\to\mathfrak{G}_{\pi}.$
-The pushforward distribution on grains, $\hat{P}_{\mathfrak{G}}:=(\ell\circ q_{\pi})_{*}\hat{P}\in\Delta(\mathfrak{G}_{\pi})$, is
-$\hat{P}_{\mathfrak{G}}(G):=\sum_{u\in(\ell\circ q_{\pi})^{-1}(G)}\hat{P}(u)\qquad(G\in\mathfrak{G}_{\pi}),$
-so that
-$\hat{P}_{\mathfrak{G}}(G_{\pi,1})=\hat{P}(1)+\hat{P}(2)+\hat{P}(3)=\frac{3}{5},\qquad\hat{P}_{\mathfrak{G}}(G_{\pi,2})=\hat{P}(4)+\hat{P}(5)+\hat{P}(6)=\frac{2}{5}.$
-4.2 Kullback–Leibler Divergence
-We quantify the information loss caused by coarse-graining via the Kullback–Leibler (KL) divergence between probability distributions due to its well-established role in measuring the divergence between probability distributions in information theory [kullbackleibler1951].
-$D_{\text{KL}}(P\parallel Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}$ (1)
-While alternative measures such as Jensen-Shannon Divergence or Wasserstein Distance may be explored in future studies, $D_{\text{KL}}$ provides a mathematically straightforward and interpretable baseline for assessing information loss in coarse evaluations.
-To illustrate the motivation behind coarse ethics, consider now that we are a small classroom teacher and give a math test to ten students. The results appear in Table 1, Figure 1.
+Given the underlying scale $(U,\leq)$, let $\pi$ be a coarse-grained partition of $(U,\leq)$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$, and let $P_{U}$ be a probability distribution on $(U,2^{U})$.
+With $q_{\pi}$ as in Definition 2, we define the coarse-grained probability distribution $P_{\pi}$ on $(I_{\pi},2^{I_{\pi}})$ as follows.
+First, for each label $i\in I_{\pi}$, set
+$P_{\pi}(\{i\}):=P_{U}(G_{\pi,i}).$
+We then extend this to arbitrary subsets $A\subseteq I_{\pi}$ by additivity, defining
+$P_{\pi}(A):=\sum_{i\in A}P_{\pi}(\{i\})=\sum_{i\in A}P_{U}(G_{\pi,i})=P_{U}\!\left(\bigcup_{i\in A}G_{\pi,i}\right).$
+Equivalently, for every $A\subseteq I_{\pi}$,
+$P_{\pi}(A)=P_{U}\bigl(q_{\pi}^{-1}(A)\bigr),$
+that is, $P_{\pi}$ is the pushforward of $P_{U}$ along $q_{\pi}$.
+Example 3 (Pushforward to a pass/fail outcome).
+As a running example, consider now that we are a small classroom teacher and give a math test to ten students. The results appear in Table 1, Figure 1.
 Table 1: Student Scores
@@
 $O:=\{s_{1},s_{2},\dots,s_{10}\}.$
-Define the score map $f:O\to U$ from Table 1 by
+Define the object-to-score map $f:O\to U$ from Table 1 by
 $f(s_{1})=5,\ f(s_{2})=3,\ f(s_{3})=3,\ f(s_{4})=3,\ f(s_{5})=10,$
@@
 $\hat{P}(u):=\frac{\bigl|\{s\in O:\ f(s)=u\}\bigr|}{|O|}\qquad(u\in U).$
-From the score list, the counts are
-$\#\{u=0\}=1,\ \#\{u=2\}=1,\ \#\{u=3\}=3,\ \#\{u=4\}=1,$
-$\#\{u=5\}=2,\ \#\{u=6\}=1,\ \#\{u=10\}=1,$
-and all other scores occur 0 times. Hence
+Hence
 $\hat{P}(0)=\frac{1}{10},\ \hat{P}(1)=0,\ \hat{P}(2)=\frac{1}{10},\ \hat{P}(3)=\frac{3}{10},\ \hat{P}(4)=\frac{1}{10},\$
@@
 $\ell(1)=\text{fail},\qquad\ell(2)=\text{pass}.$
-Define the coarse assignment map $q_{\pi}:U\to I_{\pi}$ by
+Define the score-to-category map $q_{\pi}:U\to I_{\pi}$ by
 $q_{\pi}(u)=\begin{cases}1,&u\in G_{\pi,1},\\ 2,&u\in G_{\pi,2}.\end{cases}$
@@
 $(\ell\circ q_{\pi}\circ f)(s_{10})=\text{fail}.$
-Define the pushforward distribution on labels by
-$\hat{P}_{L}:=(\ell\circ q_{\pi})_{*}\hat{P}\in\Delta(L),\qquad\hat{P}_{L}(\lambda):=\sum_{u\in(\ell\circ q_{\pi})^{-1}(\lambda)}\hat{P}(u)\quad(\lambda\in L).$
+Define the pushforward distribution on labels $\hat{P}_{L}$ on $(L,2^{L})$ by
+$\hat{P}_{L}(A):=\hat{P}\bigl((\ell\circ q_{\pi})^{-1}(A)\bigr)\qquad(A\subseteq L).$
+In particular, for a singleton $A=\{\lambda\}$ with $\lambda\in L$, we have
+$\hat{P}_{L}(\{\lambda\})=\hat{P}\bigl((\ell\circ q_{\pi})^{-1}(\{\lambda\})\bigr)=\sum_{u\in(\ell\circ q_{\pi})^{-1}(\{\lambda\})}\hat{P}(u).$
 Thus,
@@
 Equivalently, $\hat{P}_{L}(\text{fail})=\hat{P}(G_{\pi,1})$ and $\hat{P}_{L}(\text{pass})=\hat{P}(G_{\pi,2})$.
-We compare the probability distributions between the original grading system (ranging from 0 to 10) and the binary pass/fail assessment.
-More intuitively, if we inform ten students only of the pass/fail criteria and their respective pass or fail status, without disclosing their specific scores—or if we later lose the list of scores and can only recall the pass/fail criteria and each student’s pass or fail status—our task is to determine how much information has been lost.
+4.2 Kullback–Leibler Divergence
+4.2.1 Setup
+We quantify the information loss caused by coarse-graining via the Kullback–Leibler (KL) divergence between probability distributions due to its well-established role in measuring the divergence between probability distributions in information theory [kullbackleibler1951].
+$D_{\text{KL}}(P\parallel Q)=\sum_{i}P(i)\log\frac{P(i)}{Q(i)}$ (1)
+While alternative measures such as Jensen-Shannon Divergence or Wasserstein Distance may be explored in future studies, $D_{\text{KL}}$ provides a mathematically straightforward and interpretable baseline for assessing information loss in coarse evaluations.
+Using Example 3, we compare the distributions induced by the original 0–10 scoring scheme and by its binary pass/fail coarsening.
+Intuitively, if only pass/fail outcomes are retained (and the exact scores are discarded), we would like to quantify how much information is lost.
 When using $D_{\text{KL}}$ to achieve this objective, determining $P(i)$ is straightforward.
@@
 For instance, $P(0)$ is $\frac{1}{10}$, and $P(1)$ is 0 as to Table 1.
-4.3 Categorical Unification
+4.2.2 Categorical Unification
 The challenge lies in determining $Q(i)$.
@@
 At the same time, once we coarse-grain the score scale, we want the coarsened distribution to preserve the total probability mass assigned to each grain, and we would like to define a canonical representative distribution that introduces no additional structure beyond this coarse information.
-To this end, we define Categorical Unification as follows.
-Definition 5 (Categorical unification (CU) and $D_{\mathrm{KL\text{-}CU}}$).
-Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$. Let $P_{U}$ be a probability distribution on $U$.
-(i) The coarse-grained distribution induced by $\pi$ is the pushforward $P_{\pi}$ on $I_{\pi}$ defined by
-$P_{\pi}(i):=P_{U}(G_{\pi,i})=\sum_{u\in G_{\pi,i}}P_{U}(u),\qquad i\in I_{\pi}.$
-Here, the category index $i$ represents the grain $G_{\pi,i}\subseteq U$; hence $P_{\pi}(i)$ is the total probability mass assigned to that grain, obtained by summing $P_{U}(u)$ over all $u\in G_{\pi,i}$.
-(ii) The CU-unification of $P_{U}$ (through $\pi$) is the probability distribution $Q^{\mathrm{CU}}$ on $U$ defined by
-$Q^{\mathrm{CU}}(u):=\frac{P_{\pi}(i)}{|G_{\pi,i}|}\qquad\text{for }u\in G_{\pi,i}.$
-Here, $Q^{\mathrm{CU}}$ spreads the coarse-grained mass $P_{\pi}(i)$ uniformly across the grain $G_{\pi,i}$; thus each $u\in G_{\pi,i}$ receives probability $Q^{\mathrm{CU}}(u)=P_{\pi}(i)/|G_{\pi,i}|$.
+To this end, we define Categorical Unification (CU) as follows.
+Definition 5 (Categorical unification and $D_{\mathrm{KL\text{-}CU}}$).
+Given the underlying scale $(U,\leq)$, let $P_{U}$ be a probability distribution on $(U,2^{U})$.
+(i) The coarse-grained distribution induced by $\pi$ is the probability distribution $P_{\pi}$ on $I_{\pi}$ defined by
+$P_{\pi}(\{i\}):=P_{U}(G_{\pi,i})=\sum_{u\in G_{\pi,i}}P_{U}(\{u\}),\qquad i\in I_{\pi}.$
+Equivalently, in the discrete setting we write $P_{U}(u):=P_{U}(\{u\})$ and $P_{\pi}(i):=P_{\pi}(\{i\})$.
+(ii) The CU-unification of $P_{U}$ (via $\pi$) is the probability distribution $Q^{\mathrm{CU}}$ on $U$ defined by
+$Q^{\mathrm{CU}}(u):=\frac{P_{\pi}(\{i\})}{|G_{\pi,i}|}\qquad\text{for }u\in G_{\pi,i}.$
+Since $\pi$ is a partition, each $u\in U$ belongs to a unique grain $G_{\pi,i}$, so $Q^{\mathrm{CU}}$ is well-defined.
+Note that $Q^{\mathrm{CU}}$ is indeed a probability distribution on $U$: it redistributes each coarse mass $P_{\pi}(\{i\})$ uniformly over the points $u\in G_{\pi,i}$.
+In particular,
+$\sum_{u\in U}Q^{\mathrm{CU}}(u)=\sum_{i\in I_{\pi}}\sum_{u\in G_{\pi,i}}\frac{P_{\pi}(\{i\})}{|G_{\pi,i}|}=\sum_{i\in I_{\pi}}P_{\pi}(\{i\})=1.$
 (iii) The KL divergence by categorical unification is
 $D_{\mathrm{KL\text{-}CU}}(P_{U}):=D_{\mathrm{KL}}\!\left(P_{U}\,\middle\|\,Q^{\mathrm{CU}}\right),$
-where $D_{\mathrm{KL}}(P\|Q)=\sum_{u\in U}P(u)\log\!\frac{P(u)}{Q(u)}$ with the usual convention $0\log(0/q)=0$.
+where
+$D_{\mathrm{KL}}(P\|Q)=\sum_{u\in U}P(u)\log\!\frac{P(u)}{Q(u)},$
+with the convention that any term with $P(u)=0$ is taken to be 0.
+The point of Definition 5 is not to compare the original distribution directly with the coarse representation itself.
+Rather, $D_{\mathrm{KL\text{-}CU}}$ measures how far the original distribution deviates from a canonical fine-scale reconstruction induced by the coarse representation under minimal assumptions.
 The empirical rationale for this method is as follows: When students are evaluated using a coarse-grained grading system where only categorical results (e.g., pass or fail) are recorded, their exact scores become irretrievable.
@@
 Since no particular score within the pass category has more supporting evidence than another, a uniform assumption ensures neutrality and avoids introducing arbitrary weighting.
-The principle of maximum entropy provides a theoretical justification for CU.
-It states that among all possible probability distributions satisfying a given set of constraints, the one that maximizes entropy (i.e., uncertainty) is the least biased and most informative distribution [jaynes1957].
+Additionally, the principle of maximum entropy provides a theoretical justification for CU.
+It states that among all probability distributions satisfying a given set of constraints, the one that maximizes entropy (i.e., uncertainty) is the least biased and least informative beyond the imposed constraints [jaynes1957].
 In the case of pass/fail classification, however, we are not only constrained by the requirement that probabilities sum to one but also by the fact that we know the number of students in each category.
@@
 Proposition 2 (Maximum-entropy property of CU).
-Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$. Fix $P\in\Delta(U)$ and write
+Given the underlying scale $(U,\leq)$, fix $P\in\Delta(U)$ and write
 $p_{i}:=P(G_{\pi,i})=\sum_{x\in G_{\pi,i}}P(x)\qquad(i\in I_{\pi}).$
-Define the categorical unification (CU) of $P$ (relative to $\pi$) by
+Define the CU of $P$ (relative to $\pi$) by
 $P^{\mathrm{CU}}(x):=\frac{p_{i}}{|G_{\pi,i}|}\quad\text{for }x\in G_{\pi,i}.$
@@
 Then $P^{\mathrm{CU}}$ maximizes Shannon entropy $H(Q):=-\sum_{x\in U}Q(x)\log Q(x)$ over $\mathcal{C}(p,\pi)$. Moreover, if $p_{i}>0$ for all $i$, the maximizer is unique.
+The constraint set $\mathcal{C}(p,\pi)$ fixes only the coarse information, namely the total mass $p_{i}$ assigned to each grain $G_{\pi,i}$, while leaving the distribution within each grain free.
+For example, if a grain is $G_{\pi,i}=\{8,9,10\}$ and $p_{i}=0.3$, then any assignment of nonnegative probabilities $(Q(8),Q(9),Q(10))$ with $Q(8)+Q(9)+Q(10)=0.3$ is admissible, such as $(0.1,0.1,0.1)$ or $(0.29,0.01,0)$.
+Proposition 2 shows that, among all fine-grained distributions consistent with these coarse masses, the CU distribution $P^{\mathrm{CU}}$ is the least informative one in the sense of maximum Shannon entropy (in the above example, this corresponds to the uniform assignment $(0.1,0.1,0.1)$ on $G_{\pi,i}=\{8,9,10\}$): it preserves the grain-level totals $p_{i}$ and otherwise spreads mass uniformly within each grain.
+In particular, CU can be viewed as the canonical “uninformative” refinement of the coarse-grained distribution back to the original scale $U$.
 Proof of Proposition 2.
@@
 so that $\sum_{x\in G_{\pi,i}}Q_{i}(x)=1$. A direct calculation yields the entropy decomposition
-$H(Q)=H(p)+\sum_{i\in I_{\pi}}p_{i}\,H(Q_{i}),$
-where $H(p):=-\sum_{i}p_{i}\log p_{i}$ depends only on $p$ (hence is constant over $\mathcal{C}(p,\pi)$).
-Therefore, maximizing $H(Q)$ over $\mathcal{C}(p,\pi)$ reduces to maximizing each $H(Q_{i})$ over the simplex $\Delta(G_{\pi,i})$.
+$H(Q)=H(p)+\sum_{i\in I_{\pi}}p_{i}\,H(Q_{i}).$
+This identity shows that the total entropy of $Q$ consists of the entropy $H(p)$ at the coarse-grained level (which grain is selected) plus the average within-grain entropy $\sum_{i\in I_{\pi}}p_{i}H(Q_{i})$, weighted by the grain masses $p_{i}$.
+$H(p):=-\sum_{i}p_{i}\log p_{i}$ depends only on $p$ (hence is constant over $\mathcal{C}(p,\pi)$) Therefore, maximizing $H(Q)$ over $\mathcal{C}(p,\pi)$ reduces to maximizing each $H(Q_{i})$ over the simplex $\Delta(G_{\pi,i})$.
 For a fixed finite set $G_{\pi,i}$, Shannon entropy is maximized uniquely by the uniform distribution on $G_{\pi,i}$, i.e., $Q_{i}(x)=1/|G_{\pi,i}|$ for all $x\in G_{\pi,i}$.
@@
 Theorem 3 (Zero information loss under CU (finite/discrete)).
-Let $U$ be a finite set and let $\pi$ be a partition of $U$ with grain family $\mathfrak{G}_{\pi}=\{G_{\pi,i}\}_{i\in I_{\pi}}$.
-Let $P_{U}$ be a probability distribution on $U$, and define the induced coarse-grained distribution
+Given the underlying scale $(U,\leq)$, let $P_{U}$ be a probability distribution on $U$, and define the induced coarse-grained distribution
 $P_{\pi}(i):=\sum_{u\in G_{\pi,i}}P_{U}(u),\qquad i\in I_{\pi}.$
-Define the CU-unification $Q^{\mathrm{CU}}$ on $U$ by
+Define the $Q^{\mathrm{CU}}$ on $U$ by
 $Q^{\mathrm{CU}}(u):=\frac{P_{\pi}(i)}{|G_{\pi,i}|}\qquad\text{for }u\in G_{\pi,i}.$
@@
 $D_{\mathrm{KL}}\!\left(P_{U}\,\middle\|\,Q^{\mathrm{CU}}\right)=0\quad\Longleftrightarrow\quad\forall i\in I_{\pi},\ \forall u\in G_{\pi,i}:\ P_{U}(u)=\frac{P_{\pi}(i)}{|G_{\pi,i}|}.$
-Equivalently, for every $i\in I_{\pi}$ with $P_{\pi}(i)>0$, the conditional distribution of $P_{U}$ given the event $\{u\in G_{\pi,i}\}$ is uniform on $G_{\pi,i}$:
-$P_{U}(\,\cdot\,\mid G_{\pi,i})=\mathrm{Unif}(G_{\pi,i}).$
+Equivalently, for every $i\in I_{\pi}$ with $P_{\pi}(i)>0$ and every $A\subseteq G_{\pi,i}$,
+$P_{U}(A\mid G_{\pi,i})=\frac{|A|}{|G_{\pi,i}|}.$
 Proof of Theorem 3.
@@
 which is precisely $\mathrm{Unif}(G_{\pi,i})$. ∎
-4.4 Minimizing Information Loss
-We aim to make the probability distribution of the coarse-grained pass/fail evaluation as similar as possible to the original probability distribution of scores ranging from 0 to 10.
-Here, “similar” specifically refers to minimizing information loss, which is quantitatively measured by minimizing the $D_{\text{KL-CU}}$.
+This theorem characterizes the situation in which elements that are mapped to the same grain (that is, the same category) after coarsening are already assigned uniform probability mass within that grain under the pre-coarsening distribution $P_{U}$.
+4.3 Optimizing Information Loss
+4.3.1 Minimizing the KL Divergence
+When moving from simple examples such as pass/fail classification to the general design of coarse-grained evaluations, an additional question arises: how should we choose among multiple possible coarse-grainings?
+Once admissible coarse-grainings have been restricted to CGPs satisfying the structural requirements of CE, the remaining question is how to compare multiple admissible partitions.
+The role of $D_{\mathrm{{KL\text{-}CU}}}$ in this section is to provide one such comparative criterion.
+As the zero-loss theorem for $D_{\mathrm{KL\text{-}CU}}$ shows, making information loss vanish in the sense of this measure requires that the original fine-grained distribution already be uniform within each grain.
+In practice, however, the criteria for coarsening an evaluation are rarely chosen so as to align with the original distribution in this way.
+For example, in grading a mathematics test, a teacher would not normally create a category simply because the numbers of students scoring 61, 62, and 63 happen to be equal.
+The zero-loss theorem therefore should not be understood as providing a practical guideline for optimizing categorization.
+Rather, it shows that practical categorization is usually guided by considerations other than the complete elimination of information loss in the sense of $D_{\mathrm{KL\text{-}CU}}$.
+One example in which zero information loss cannot be expected is when the number of categories is fixed in advance—for instance, when outcomes must be classified only as pass or fail.
+One possible idea, then, is to give up on achieving zero information loss and instead set the threshold so as to minimize information loss.
 Suppose we need to divide those ten students (Table 1) into pass and fail groups, but we can determine the passing threshold $(T)$.
 What score should be used as $T$?
-A seemingly natural approach is to use the average score.
-Since the mean score in this test is 4.1, setting the passing threshold at $T=4$ may seem reasonable.
-If we classify students scoring below 4 as failing and those scoring 4 or above as passing, we obtain a $D_{\text{KL-CU}}$ of approximately 0.524, which is relatively low.
-However, this is not the minimum.
+In this setting, when the number of grains is fixed, minimizing information loss provides a natural yet straightforward way to formulate the objective function.
 The lowest $D_{\text{KL-CU}}$ is achieved when the passing threshold is set at $T=7$, resulting in a value of 0.381 (Table 4, Figure 2).
@@
 Figure 2: Graphing of Table 4
-If an AI teacher performed this calculation and told a human teacher, “Based on this test, only $s_{5}$ should pass”, the human teacher might find it counterintuitive.
-However, its interpretation is context-dependent.
-If the number of passing students is fixed (e.g., a predetermined quota of three students), minimizing $D_{\text{KL-CU}}$ may not always be the correct objective.
-In such cases, the goal is not to preserve the original probability distribution but rather to adhere to external constraints, such as ensuring that exactly three students pass, regardless of their relative performance.
-This suggests that setting a fixed pass quota can significantly distort the original probability distribution.
-Many educators may have encountered situations in which students had widely varying abilities or had to pass students whose performance was lower than they would ideally consider acceptable.
-This raises a critical question: If the information loss is excessively large, should AI provide an advisory to the teacher?
-To address this, it may be beneficial to equip the AI teacher with a feature that advises, “The school has mandated that three students must pass this test.
-Do you believe this number is truly appropriate?” This would allow for human oversight in cases where strict adherence to an imposed quota might compromise fairness or accuracy.
-In some contexts, setting the threshold at $T=7$, where only $s_{5}$ passes, may actually align with intuition.
-Consider a scenario in which the goal is to divide the class into two distinct groups: one for mathematically advanced students (e.g., an honors class) and another for those who need more support.
-The original score distribution implies that only one student significantly outperforms the others in mathematics; accordingly, $T=7$ identifies a single clear outlier ($s_{5}$).
-Thus, setting $T=7$ produces a classification that best preserves the structure of the original probability distribution while maintaining a meaningful distinction between high and low performers.
-This analysis highlights how the calculation of $D_{\text{KL-CU}}$ can serve as a bridge between AI-driven decision-making and human intuition, offering a foundational example for both XAI and CE.
+However, reducing the coarse-graining design problem to mere information-loss minimization is not always adequate.
+In many settings (e.g., pass/fail decisions), the relevant objective is not to reproduce the original score distribution, but rather to minimize decision costs (such as false positives and false negatives) or to optimize expected utility.
+Consider again the arithmetic test example above.
+If the pass threshold $T$ is chosen so as to minimize $D_{\mathrm{KL\text{-}CU}}$, the optimum may be a threshold such as $T=7$, under which only the student $s_{5}$ passes.
+This is reasonable from the viewpoint of distributional fidelity, because it preserves, within the coarse pass/fail representation, the fact that $s_{5}$ stands out as an exceptionally strong performer in this class of ten students.
+Yet such an information-loss-minimization principle is justified only when the teacher’s primary aim is to preserve information about the original score distribution.
+In educational practice, pass/fail evaluations are often tied instead to an operational criterion, such as whether a student can keep up with the next course module.
+For example, if scoring at least 6 points is sufficient to follow the next class, the teacher may reasonably wish to set the pass threshold at $T=6$.
+In that case, selecting $T=7$ by minimizing $D_{\mathrm{KL\text{-}CU}}$ may still be meaningful from the viewpoint of distributional fidelity, but it fails to reflect the actual decision objective and therefore becomes an inappropriate design principle for the coarse-graining.
+4.3.2 An Optimization View of Coarse-Grained Design
+Therefore, KL-based information-loss minimization should be understood as a principled baseline objective, one that may need to be supplemented or replaced when decision-theoretic or normative constraints are dominant.
+Formally, given a finite totally ordered set $(U,\leq)$ and a distribution $P$ on $U$, we consider the optimization problem
+$\min_{\pi\in\mathcal{P}_{\mathrm{int}}(U,\leq)}\;\;D_{\mathrm{KL}}\!\left(P\,\middle\|\,Q^{\mathrm{CU}}_{\pi}\right)\;+\;\lambda\,\Omega(\pi),$ (2)
+where $\mathcal{P}_{\mathrm{int}}(U,\leq)$ denotes the collection of interval coarse-grained partitions, $Q^{\mathrm{CU}}_{\pi}$ is the CU-unification associated with $\pi$, $\Omega(\pi)$ is a complexity penalty (e.g., the number of grains $m(\pi)$ or a description-length proxy), and $\lambda\geq 0$ controls the trade-off between informational fidelity and simplicity.
+In this formulation, coarse-graining is not merely an evaluative act, but also the choice of a representation guided by optimization, and one that can be adapted to domain-specific constraints and interpretability demands.
+The case $\lambda=0$ corresponds to the limiting situation in which minimizing $D_{\mathrm{KL\text{-}CU}}$ is itself the sole objective.
+In that case, the finest partition is trivially optimal, since no penalty is imposed on excessive refinement.
+The present framework is intended primarily for the nontrivial case $\lambda>0$, where the design of coarse-graining requires a trade-off between informational fidelity and coarsening cost.
+From this perspective, the role of $\lambda$ is not merely technical, but expresses the practical fact that coarse-grained evaluation is introduced precisely because interpretability, simplicity, and other costs of fine-grained representation cannot be ignored.
+We leave a systematic study of (2), including computational aspects and constraint variants, to future work.
 5 Discussion
-5.1 Mathematical Foundation of Coarse Ethics
-CGPs have demonstrated that the rationality of coarse evaluations, as argued by CE, can be rigorously analyzed.
-The coarse evaluations we use in daily life—for instance, the assessment that “Olivia received an excellent grade on her mathematics test”—exhibit a structured mathematical form within CGPs.
-The fact that an evaluation is coarse—for example, that Olivia’s specific score difference from other students who also received an “excellent” rating is intentionally disregarded—does not imply that the evaluation is ambiguous.
-Here, we can clearly distinguish between coarseness and ambiguity in assessment.
-Traditional ethical theories, particularly consequentialist (e.g., felicific calculus by Bentham [bentham1789introduction]) and virtue-based approaches (e.g., Stoicism), often emphasize continuity in moral evaluation, where actions that align more closely with a given ideal are ranked proportionally higher.
-However, CE suggests that even when individuals differ significantly in their alignment with an ideal, they may still be evaluated equivalently [izumoweng2022].
-This aligns with Immanuel Kant’s criticism of excessive ethical precision, particularly his rejection of casuistry—the attempt to make moral laws overly specific through case-by-case analysis [kant2017].
-A simple example illustrates this point: Suppose a student receives a failing grade of 30 points on an essay from a human teacher.
-Dissatisfied, they submit the same essay to an AI teacher for evaluation, which assigns them a score of 30.1 while their peer, who was also given 30 by the human teacher, receives 29.8 from the AI.
-If the student then argues that the AI’s refined scoring justifies a reconsideration of their grade, such a claim would be unreasonable.
-From the teacher’s perspective, both evaluations reflect insufficient performance, and fixating on a marginal difference of 0.3 points does not meaningfully change the outcome.
-5.2 Tradeoff Between Evaluation Efficiency and Information Retention
-As demonstrated in this study, when $Q(i)$ in $D_{\text{KL}}$ is processed through categorical unification, information is preserved only in specific cases.
-Our findings suggest a potential trade-off between evaluation efficiency and information retention.
-While this study does not provide a formal proof, we hypothesize that as evaluation processes prioritize simplicity and accessibility, the loss of fine-grained details becomes inevitable.
-Further research could investigate whether there exists a Pareto-optimal balance between evaluation efficiency and retained information in AI-driven decision-making.
-Assuming a uniform distribution within a category is the most reasonable default when no additional information is available.
-When only a pass label is recorded, any distinction between different passing scores is inherently erased, making a categorically uniform distribution the simplest assumption that reflects this lack of distinguishing data.
-While alternative assumptions, such as normal or skewed distributions, might seem more refined, they require external data that is often unavailable in real-world grading systems.
-If historical trends suggest a particular distribution (e.g., most passing students score near the middle of the range), Bayesian updates could provide a better-informed probability model.
-However, in the absence of such priors, assuming a categorically uniform distribution remains the least biased approach.
-5.3 Application to Explainable AI
-We apply the above discussion to the challenges identified in Section 1.2.
-One key application is in refining AI-driven safety advisories.
-For example, in autonomous driving systems, the coarse-graining of safety warnings must not result in the loss of critical risk information.
-Instead of issuing a generic warning such as “Your driving is dangerous”, CGPs can guide AI systems to adaptively determine the level of warning granularity, ensuring that important details about immediate risks (e.g., speed excess relative to road conditions or proximity to nearby obstacles) are retained within the coarsened message.
-This approach allows for a structured balance between reducing cognitive overload and preserving necessary safety information.
-Additionally, CGPs offer a refined approach to addressing the limitations of the three-tiered explanatory model in CE.
-Rather than strictly categorizing explanations into those for experts, general users, and individuals with cognitive difficulties, CGPs allow for dynamic adaptation of explanation granularity by integrating user profiling mechanisms.
-In a medical diagnostic AI, for example, an expert might receive a detailed probability breakdown of differential diagnoses, while a general patient would be presented with a high-level summary emphasizing key risks and actionable next steps.
-This ensures that explanations remain accessible without sacrificing necessary details for informed decision-making.
-This personalized granularity adjustment can improve the effectiveness of AI explanations in domains such as medical diagnostics, legal decision-making, and financial advising, where different users require distinct levels of detail in their assessments.
+We show that the core ideas of CE can be given a precise mathematical formulation.
+Coarse evaluations in everyday life—for example, the statement that “Olivia received an excellent grade on her mathematics test”—exhibit a clear mathematical structure within the framework of coarse-grained partitions.
+The fact that an evaluation is coarse—for instance, that Olivia’s exact score difference from other students who also received an “excellent” rating is intentionally ignored—does not mean that the evaluation is ambiguous.
+In this sense, CGPs make it possible to distinguish clearly between coarseness and ambiguity in assessment.
+Therefore, CE’s attempt to distinguish coarse evaluations from arbitrary or haphazard evaluations can be justified mathematically.
+CE has proposed, as a criterion for comparing fine-grained and coarse-grained evaluations, that a coarse-grained evaluation should not reverse the value ordering embodied in the fine-grained one.
+However, as the previous literature itself acknowledges, multiple coarse-grained evaluations can satisfy this condition for the same fine-grained evaluation.
+The problem, therefore, is that this criterion alone cannot sufficiently determine how the coarsening should be carried out.
+In response, we have proposed a new criterion based on comparing the probability distributions associated with the fine-grained and coarse-grained evaluations, and on minimizing the sum of information loss and coarsening cost.
+When the fine-grained evaluation is given as a finite discrete totally ordered set, a coarse-grained evaluation constructed within the framework of CGPs allows the corresponding probability distributions and information loss to be defined and computed rigorously.
+Moreover, the central result of this paper—the zero-loss theorem for $D_{\mathrm{KL\text{-}CU}}$—shows that coarsening yields zero information loss in the sense of this measure only when the original fine-grained distribution is already uniform within each grain.
+Hence, coarse-grained evaluations with zero information loss, in the sense of $D_{\mathrm{KL\text{-}CU}}$, occur only in highly exceptional cases in actual evaluative practice.
+These observations are also relevant to explainable AI.
+In many AI systems, the model first produces a fine-grained internal evaluation of its target.
+This is true, for example, in AI-assisted driving systems that compute detailed risk scores, in medical AI that estimates diagnostic likelihoods, and in hiring or interview AI that generates fine-grained personnel assessments.
+From this perspective, one important problem in XAI is how such fine-grained internal evaluations should be transformed into coarser outputs that are intelligible and usable for human agents.
+A simple example is AI-assisted driving.
+Suppose that the system internally represents driving risk on a fine-grained scale such as $U=\{0,1,\dots,100\}$, while the human-facing interface must display only a small number of warnings, such as safe, caution, and danger.
+In this case, a CGP represents the transformation from the fine-grained risk scale to these coarse warning categories.
+The corresponding quantity $D_{\mathrm{KL\text{-}CU}}$ then measures how far the original fine-grained risk distribution deviates from the canonical fine-scale reconstruction induced by that coarse warning system under minimal assumptions.
+This makes the design problem more precise.
+The issue is not only whether the warning system remains intelligible to the driver, but also how much fine-grained risk information is lost when the internal evaluation is compressed into a small number of human-interpretable categories.
+In such cases, the framework proposed in this paper becomes directly relevant: a principled warning system should balance informational fidelity against the coarsening cost associated with human cognitive burden, instruction complexity, and response usability.
 6 Conclusion
-This study has established a rigorous mathematical foundation for Coarse Ethics (CE).
-By introducing the framework of CGPs, we demonstrated that coarse evaluations can be systematically analyzed rather than being treated as vague or arbitrary assessments.
-A coarse set is defined by constructing a coarse-grained partition over a totally ordered set, and its mathematical properties have been formalized through the definitions, propositions, and theorems in this paper.
-The key conclusions drawn from this study are as follows:
-  1. 1.
-    Coarse evaluation methods, as described in CE, can be formulated with mathematical precision, providing a structured approach to assessments.
-  2. 2.
-    Different coarse evaluation methods can be analyzed within the framework of CGPs.
-In particular, each evaluation method can be represented as a probability distribution, allowing for probabilistic and statistical comparison.
-  3. 3.
-    A quantitative approach to evaluating information loss in coarse assessments is provided by Kullback-Leibler Divergence, which measures the deviation of a coarse-grained distribution from the original evaluation.
-As AI systems become more complex, the need for interpretable and adaptable explanations grows.
-Current XAI techniques, such as LIME and SHAP, often provide explanations at a fixed granularity, which may not align with user-specific cognitive requirements.
-CGPs offer a structured way to dynamically adjust the level of detail in AI explanations, ensuring that different users receive information at an appropriate level of abstraction.
-This approach could enhance AI-assisted decision-making in critical fields such as medicine, finance, and law, where balancing interpretability and complexity is essential.
-This study presents an initial exploration of CGPs, but several limitations remain.
-Our analysis is restricted to totally ordered sets and assumes a covering condition, meaning that every element of the original set is included within the coarse-grained structure.
-This differs from previous research, which allows for the relaxation of coverage constraints.
-In other words, our approach focuses on cases where a coarse evaluation fully encompasses the set of evaluated entities.
-Additionally, our discussion has been primarily framed within the context of educational assessments, particularly school testing.
-Future research should investigate the implications of relaxing the covering condition and explore applications in broader ethical and legal decision-making frameworks.
+This paper has provided a rigorous mathematical framework for analyzing coarse-grained evaluation in the spirit of Coarse Ethics (CE).
+By introducing coarse-grained partitions (CGPs) on finite discrete totally ordered sets, we showed that coarse evaluations need not be treated as vague or arbitrary, but can instead be represented and analyzed in a precise mathematical form.
+Our main contribution is threefold.
+First, we formalized coarse-grained evaluation as a mapping from a fine-grained score scale to a coarser categorical scale.
+Second, we introduced categorical unification (CU) and thereby defined a KL-based measure, $D_{\mathrm{KL\text{-}CU}}$, of how far the original fine-grained distribution departs from a canonical fine-scale reconstruction induced by the coarse representation under minimal assumptions.
+Third, we proved that this divergence vanishes only under a highly restrictive condition: within each grain, the original fine-grained distribution must already be uniform.
+Hence, zero information loss in the sense of $D_{\mathrm{KL\text{-}CU}}$ is not a realistic benchmark for ordinary evaluative practice, but rather a highly exceptional limiting case.
+These results also lead naturally to the problem of optimization.
+Since multiple coarse-grained evaluations may be compatible with the same fine-grained evaluation, the issue is not merely whether coarse-graining is possible, but how it should be designed.
+Our analysis suggests that minimizing KL-based information loss provides a principled baseline objective, although in real settings this objective may need to be supplemented or replaced by operational, normative, or decision-theoretic considerations.
+In this sense, the framework developed here should be understood not as a universal rule for categorization, but as a formal basis for comparing alternative design objectives.
+Finally, the framework is also relevant to explainable AI.
+In AI-assisted settings, a system may hold a fine-grained internal evaluation while needing to present a coarser representation that is intelligible to human users.
+This can be understood as a transformation from a fine-grained evaluation to a coarse-grained one, and our framework provides a natural way to study the trade-off between preserving critical information and reducing coarsening costs such as cognitive burden and explanatory complexity.
+Future work should extend the present analysis beyond finite discrete ordered settings and investigate richer optimization problems under domain-specific constraints.
 Appendix A. Python Code for Calculating $D_{\text{KL-CU}}$